### v1

# The inducibility of Turán graphs

Abstract
Let $I(F,n)$ denote the maximum number of induced copies of a graph $F$ in an $n$-vertex graph. The inducibility of $F$, defined as $i(F)=\lim_{n\to\infty}{I(F,n)}/{\tbinom{n}{v(F)}}$, is a central problem in extremal graph theory. In this work, we investigate the inducibility of Turán graphs $F$. This topic has been extensively studied in the literature, including works of Pippenger–Golumbic [27], Brown–Sidorenko [7], Bollobás–Egawa–Harris–Jin [5], Mubayi, Reiher, and the first author [23], and Yuster [32]. Broadly speaking, these results resolve or asymptotically resolve the problem when the part sizes of $F$ are either sufficiently large or sufficiently small (at most four).
We complete this picture by proving that for every Turán graph $F$ and sufficiently large $n$, the value $I(F,n)$ is attained uniquely by the $m$-partite Turán graph on $n$ vertices, where $m$ is given explicitly in terms of the number of parts and vertices of $F$. This confirms a conjecture of Bollobás–Egawa–Harris–Jin [5] from 1995, and we also establish the corresponding stability theorem. Moreover, we prove an asymptotic analogue for $I_{k+1}(F,n)$, the maximum number of induced copies of $F$ in an $n$-vertex $K_{k+1}$-free graph, thereby completely resolving a recent problem of Yuster [32]. Finally, our results extend to a broader class of complete multipartite graphs in which the largest and smallest part sizes differ by at most on the order of the square root of the smallest part size.
1 Introduction
A fundamental problem in extremal graph theory is to determine the maximum number of induced copies of a given graph $F$ among all $n$-vertex graphs. Formally, given two graphs $F$ and $G$, let $I(F,G)$ denote the number of induced copies of $F$ in $G$, that is, the number of subsets $S\subseteq V(G)$ of size $v(F)$ such that the induced subgraph $G[S]$ is isomorphic to $F$. Here, $v(F)$ denotes the number of vertices of $F$. For every positive integer $n$, let
$I(F,n)\coloneqq\max\left\{I(F,G)\colon v(G)=n\right\}.$
The inducibility of $F$ is then defined as
$i(F)\coloneqq\lim_{n\to\infty}{I(F,n)}/{\tbinom{n}{v(F)}}$.
A systematic study of the inducibility problem for graphs was initiated in a foundational work of Pippenger–Golumbic [27], in which they established several general properties of inducibility and determined the inducibility of complete bipartite graphs with part sizes differing by at most one.
Determining $I(F,n)$ (or even $i(F)$) is rather difficult in general. For example, the inducibility of the path on four vertices remains wide open (see, e.g., [10, 9]). For small graphs, bounds on the inducibility of graphs on $4$, 55, and $6$ vertices have been obtained in several works, including Exoo [10], Hirst [18], Even-Zohar–Linial [9], Pikhurko–Sliačan–Tyros [26], and Bodnár et al. [4], some of which employ the computer-assisted flag algebra machinery of Razborov [28]. A particularly interesting result obtained by Balogh–Hu–Lidickỳ–Pfender [2] is the determination of the inducibility of the $5$-cycle $C_{5}$, which confirms a special case of an old conjecture of Pippenger–Golumbic [27] on the inducibility of cycles. This conjecture remains open for longer cycles, and improved upper bounds were obtained in [17, 19].
For general graphs, results of Yuster [31] and Fox–Huang–Lee [11] imply that for almost all graphs $F$,
$i(F)=\frac{v(F)!}{v(F)^{v(F)}-v(F)},$
where the lower bound construction arises from nested blowups of $F$ itself. Answering a question of Bollobás–Egawa–Harris–Jin [5] asymptotically, Hatami–Hirst–Norine [15] established that if $F$ is a sufficiently large balanced blowup of some graph $K$, then the extremal graph for $I(F,n)$ is essentially a blowup of $K$. Extending the results [24, 20, 12] on the edge-statistic conjecture of Alon–Hefetz–Krivelevich–Tyomkyn [1], Ueltzen [30] recently classified all graphs with high inducibility.
In this work, we focus on the case where $F$ is a complete multipartite graph, and more specifically, an almost balanced complete multipartite graph. This class of graphs has already been studied extensively since the work of Pippenger–Golumbic [27]. Throughout this work, for positive integers $a_{1},\ldots,a_{r}$, let $K_{a_{1},\ldots,a_{r}}$ denote the complete $r$-partite graph with part sizes $a_{1},\ldots,a_{r}$. For positive integers $\ell\geq r$, the Turán graph $T_{r}(\ell)$ denotes the complete $r$-partite graph on $\ell$ vertices in which the largest and smallest part sizes differ by at most one. The cases $\ell=r$ and $r=1$ correspond to the complete graph on $r$ vertices and the empty graph on $\ell$ vertices, respectively; both are trivial cases in the inducibility problem. So, we assume for the remainder of this work that $\ell\geq r+1$ and $r\geq 2$.
Results of Pippenger–Golumbic [27] (also Bollobás–Nara–Tachibana [6]) show that for bipartite Turán graphs $F$, the value of $I(F,n)$ is attained by the bipartite Turán graph $T_{2}(n)$. Using Zykov symmetrization [33] together with additional arguments, Brown–Sidorenko [7] showed that for every complete multipartite graph $F$, an extremal graph for $I(F,n)$ can always be chosen from the class of complete multipartite graphs; moreover, if $F$ is complete bipartite, then the number of parts can be taken to be at most two. Nevertheless, determining the number of parts and the ratios of part sizes in the extremal construction for a complete multipartite graph $F$ remains difficult in general. Consequently, the inducibility of complete multipartite graphs is still largely open.
The family of Turán graphs $T_{r}(\ell)$ has received substantial attention in the literature. Refining an asymptotic result of Brown–Sidorenko [7] for the balanced complete $r$-partite graph $K_{r}(t):=T_{r}(rt)$, Bollobás–Egawa–Harris–Jin [5] proved that when $t\geq(1+o(1))\ln r$, the $r$-partite Turán graph $T_{r}(n)$ is the unique extremal graph, whereas this is not the case if $t<\frac{\ln(r+1)}{r\ln(1+1/r)}$. In the last section of [5], they remarked that (with $f(n,K_{r}(t))$ below corresponding to $I(K_{r}(t),n)$ in our notation):
One may also venture the conjecture that for every pair $(r,t)$, $r\geq 4$, $t\geq 2$, if $n$ is sufficiently large then for some $s\geq 0$, $T_{r+s}(n)$ is the unique extremal graph for $f(n,K_{r}(t))$.
This conjecture has been open for three decades. There has been recent progress on the case when $F$ is a Turán graph with each part of small size. Mubayi, Reiher, and the first author [23, Theorem 1.13] determined the inducibility when $F$ is a Turán graph in which all but one part have size one. This result was extended very recently by Yuster [32, Theorem 1.6], who determined the inducibility when $F$ is an $r$-partite Turán graph with at most $3r+1$ vertices (in particular, all parts have size at most four). He also remarked that it would be highly interesting to determine whether the bound $3r+1$ can be removed, and posed a specific problem concerning the inducibility of Turán graphs in clique-free graphs (see Problem 1.2 for details).
The main results of this work confirm the conjecture of Bollobás–Egawa–Harris–Jin [5] and resolves the above problem of Yuster [32] in a stronger form.
1.1 Inducibility of almost balanced graphs
Let $F=K_{a_{1},\ldots,a_{r}}$ be the complete $r$-partite graph with part sizes $a_{1},\ldots,a_{r}$, and assume that $a_{1}\geq\cdots\geq a_{r}\geq 1$. Let $\ell\coloneqq a_{1}+\cdots+a_{r}$ denote the number of vertices of $F$. We say that $F$ is almost balanced if it is not the complete graph and
$\binom{a_{1}-a_{r}}{2}<a_{r},$ (1)
or, equivalently, $a_{1}<a_{r}+\tfrac{1}{2}\left(1+\sqrt{8a_{r}+1}\right)$. It is key to observe that for $\ell\geq r+1$, every Turán graph $T_{r}(\ell)$ is almost balanced. Define the constant (depending only on $F$)
$\kappa_{F}\coloneqq\frac{\binom{\ell}{a_{1},\ldots,a_{r}}}{\mathrm{sym}(a_{1},\ldots,a_{r})},$ (2)
where $\mathrm{sym}(a_{1},\ldots,a_{r})$ denotes the number of distinct permutations of the multiset $\{\hskip-2.5pt\{\hskip 1.00006pta_{1},\ldots,a_{r}\hskip 1.00006pt\}\hskip-2.5pt\}$. Let $m_{r,\ell}$ denote the unique [^1] The uniqueness will be proved in Lemma 3.2. integer that maximizes the discrete function $f\colon[r,\infty)\to\mathbb{R}$ defined by
$f(k)\coloneqq\frac{(k-1)_{r-1}}{k^{\ell-1}}=\frac{(k-1)\cdots(k-r+1)}{k^{\ell-1}},\quad\text{for every integer $k\in[r,\infty)$}.$ (3)
The main result of this subsection is the following theorem, which confirms the conjecture of Bollobás–Egawa–Harris–Jin [5] in a more general setting.
Theorem 1.1.
Suppose that $F$ is an almost balanced complete $r$-partite graph on $\ell\geq r+1$ vertices, and let $m=m_{r,\ell}$. Then
$i(F)=\kappa_{F}\cdot\frac{(m-1)_{r-1}}{m^{\ell-1}}.$
Moreover, there exists a constant $N_{F}$ such that for all $n\geq N_{F}$, the $m$-partite Turán graph $T_{m}(n)$ is the unique extremal graph for $I(F,n)$.
In the balanced case $F=K_{r}(t)$ (i.e., every part of $F$ has the same size $t$), we are able to determine the exact value of $I(F,n)$ for every positive integer $n$ (see Theorem 5.3).
It is clear that $I(F,G)=I(\bar{F},\bar{G})$, where $\bar{F}$ and $\bar{G}$ denote the complements of $F$ and $G$, respectively. Thus $I(F,n)=I(\bar{F},n)$, and hence, Theorem 1.1 also determines the inducibility of graphs that are vertex-disjoint unions of cliques, in which the smallest clique size $a_{r}$ and the largest clique size $a_{1}$ satisfy (1).
1.2 Inducibility of almost balanced graphs in $H$-free graphs
Given a graph $H$, we say that a graph $G$ is $H$-free if it does not contain $H$ as a (not necessarily induced) subgraph. Let $I_{H}(F,n)$ denote the maximum number of induced copies of $F$ in an $H$-free $n$-vertex graph, that is,
$I_{H}(F,n)\coloneqq\max\big\{I(F,G)\colon\text{$v(G)=n$ and $G$ is $H$-free}\big\}.$
The $H$-free inducibility of $F$ is then defined as
$i_{H}(F)\coloneqq\lim_{n\to\infty}{I_{H}(F,n)}/{\tbinom{n}{v(F)}}$.
When $H=K_{k+1}$ is the complete graph on $k+1$ vertices, we simply write $I_{k+1}(F,n)$ and $i_{k+1}(F)$ instead of $I_{K_{k+1}}(F,n)$ and $i_{K_{k+1}}(F)$.
A classical example is the Erdős Pentagon Problem [8], which asks for the determination of $I_{3}(C_{5},n)$, that is, the maximum number of (induced) copies of $C_{5}$ in an $n$-vertex $K_{3}$-free graph. This problem was solved independently by Grzesik [14] and Hatami–Hladký–Kráľ–Norine–Razborov [16] for large $n$, and subsequently by Lidický–Pfender [21] for all $n$.
Very recently, Yuster [32] initiated a systematic study of inducibility problem in $H$-free graphs. Among many other results, he determined (see [32, Theorem 1.6]), for all $k\geq r$, the value of $i_{k+1}(F)$ when $F$ is an $r$-partite Turán graph on at most $3r+1$ vertices. He also remarked that it would be highly interesting to determine whether the analogous result holds for all $r$-partite Turán graphs, that is, without the restriction on the number of vertices.
Problem 1.2 ([32, Problem 1.8]).
Is it true that for all $2\leq r<\ell$, there exists $t=t(r,\ell)$ such that the following holds? Let $F$ be the $r$-partite Turán graph on $\ell$ vertices. For all $k\leq t$, $i_{k+1}(F)$ is attained asymptotically by the $k$-partite Turán graphs, and for all $k\geq t+1$, $i_{k}(F)$, and hence also $i(F)$, is attained asymptotically by the $t$-partite Turán graphs.
In the following theorem, we completely resolve Yuster’s problem (in fact, for a broader family) and determine the value of $t=t(r,\ell)$, namely the integer $m_{r,\ell}$ defined in the previous subsection.
Theorem 1.3.
Suppose that $F$ is an almost balanced complete $r$-partite graph on $\ell\geq r+1$ vertices, and let $m=m_{r,\ell}$. Then
$i_{k+1}(F)=\begin{cases}\kappa_{F}\cdot\frac{(k-1)_{r-1}}{k^{\ell-1}},&\quad\text{if}\quad k\in[r,m-1],\\[3.00003pt] \kappa_{F}\cdot\frac{(m-1)_{r-1}}{m^{\ell-1}},&\quad\text{if}\quad k\in[m,\infty).\end{cases}$
1.3 Perfect stability
An interesting phenomenon, and also a very useful tool in extremal combinatorics, is stability, introduced in the seminal work of Simonovits [29]. Partly inspired by the work of Norin–Yepremyan [25], a general framework for establishing a strong form of stability in certain graph extremal problems that can be solved using the Zykov symmetrization was developed recently by Liu–Pikhurko–Sharifzadeh–Staden [22]. This framework will be used to prove the second assertion (i.e. the exact result) of Theorem 1.1.
For two graphs $G$ and $H$ with the same number of vertices, the edit distance $\mathrm{edit}(G,H)$ between $G$ and $H$ is the minimum number of edges one needs to add or remove from $G$ to make it isomorphic to $H$.
Definition 1.4 ([22, Definition 2]).
The inducibility problem for a complete multipartite graph $F$ is perfectly stable if there exists a constant $C>0$ such that for every $n$-vertex graph $G$ with $n\geq C$, there is an $n$-vertex complete multipartite graph $H$ satisfying
$\mathrm{edit}(G,H)\leq C\cdot\frac{I(F,n)-I(F,G)}{\tbinom{n}{v(F)}}\binom{n}{2}$.
In particular, perfect stability of $F$ implies that for all sufficiently large $n$, every extremal graph for the inducibility problem $I(F,n)$ is complete multipartite.
By the result of Brown and Sidorenko [7, Proposition 1], the inducibility problem for complete multipartite graphs can be solved using Zykov symmetrization. Thus, the framework from [22] can be applied to this problem. Indeed, as applications, Liu–Pikhurko–Sharifzadeh–Staden [22] established perfect stability for the inducibility problem of $F$ when $F$ is a complete bipartite graph, a complete $r$-partite graph with each part of size $t>1+\ln r$, and for small cases such as $K_{2,1,1,1}$ and $K_{3,1,1}$. They further conjectured [22, Conjecture 1] that the inducibility problem is perfectly stable for every complete multipartite graph $F$.
We confirm their conjecture for all almost balanced complete multipartite graphs.
Theorem 1.5.
Suppose that $F$ is an almost balanced complete multipartite graph. Then the inducibility problem $I(F,n)$ is perfectly stable.
Organization: In Section 2, we present some necessary definitions and preliminary results. In Section 3, we determine the values of $i(F)$ and $i_{k+1}(F)$ for almost balanced complete multipartite graphs $F$. For technical reasons, the proofs are divided into two parts: first, we handle the case $\ell\leq 2r-1$ (Subsection 3.2), and then the case $\ell\geq 2r$ (Subsection 3.3). In Section 4, we present the proof of Theorem 1.5 (see Subsections 4.2 and 4.3). In Section 5, we determine the unique extremal construction for $I(F,n)$ for sufficiently large $n$, thereby completing the proof of Theorem 1.1. The last section (Section 6) contains some concluding remarks.
2 Preliminaries
In this section, we introduce some necessary definitions and preliminary results.
For a positive integer $k$, let $k]\coloneqq\{1,\ldots,k\}$. We denote by $\mathbb{N}$ the set of natural numbers $\{0,1,2,\dots\}$, and by $\mathbb{N}_{+}$ the set of positive integers $\{1,2,3,\dots\}$. Given a (possibly infinite) set $S\subseteq\mathbb{N}$, let $(S)_{r}$ denote the collection of ordered $r$-tuples with pairwise distinct entries, that is,
$(S)_{r}\coloneqq\big\{(x_{1},\ldots,x_{r})\in S^{r}\colon\text{$x_{i}\neq x_{j}$ for all distinct $i,j\in[r]$}\big\}.$
In particular,
$(\mathbb{N})_{r}$ $\coloneqq\big\{(x_{1},\ldots,x_{r})\in\mathbb{N}^{r}\colon\text{$x_{i}\neq x_{j}$ for all distinct $i,j\in[r]$}\big\},$
$(\mathbb{N}_{+})_{r}$ $\coloneqq\big\{(x_{1},\ldots,x_{r})\in\mathbb{N}_{+}^{r}\colon\text{$x_{i}\neq x_{j}$ for all distinct $i,j\in[r]$}\big\}.$
For a complete multipartite graph $F$, let $\mathrm{sc}(F)$ denote the number of parts of $F$ that have size exactly one, that is, the number of singleton parts. We have the following simple fact for counting the number of induced copies of $F$ in a complete multipartite graph.
Fact 2.1.
Let $a_{1}\geq\cdots\geq a_{r}\geq 1$ be integers, and let $F=K_{a_{1},\ldots,a_{r}}$ be the complete $r$-partite graph with part sizes $a_{1},\ldots,a_{r}$. Suppose that $G$ is a complete multipartite graph whose non-singleton part sizes are $b_{1},\ldots,b_{k}$. Then
$I(F,G)=\frac{1}{\mathrm{sym}(a_{1},\ldots,a_{r})}\sum_{i=0}^{\mathrm{sc}(F)}i!\binom{\mathrm{sc}(F)}{i}\binom{\mathrm{sc}(G)}{i}\sum_{(i_{1},\ldots,i_{r-i})\in([k])_{r-i}}~\prod_{j=1}^{r-i}\binom{b_{i_{j}}}{a_{j}}$.
In particular, if $\mathrm{sc}(F)=0$, then
$I(F,G)=\frac{1}{\mathrm{sym}(a_{1},\ldots,a_{r})}\sum_{(i_{1},\ldots,i_{r})\in([k])_{r}}~\prod_{j=1}^{r}\binom{b_{i_{j}}}{a_{j}}$. (4)
A foundation of our approach is the following result of Brown–Sidorenko [7].
Lemma 2.2 ([7, Proposition 1]).
Suppose that $F$ is a complete multipartite graph with at least two parts. Then for every $n\geq 1$, there exists a complete multipartite graph $G$ on $n$ vertices such that $I(F,G)=I(F,n)$.
Roughly speaking, the main idea in the proof of Lemma 2.2 is to show that Zykov symmetrization does not decrease the number of induced copies of $F$. Thus, after finitely many symmetrizations, one always ends up with a complete multipartite graph. Since Zykov symmetrization also preserves $K_{k+1}$-freeness, the proof of Lemma 2.2 implies that (as observed in [32]), when $F$ is complete multipartite, there exists an extremal construction for the inducibility problem $I_{k+1}(F,n)$ that is complete multipartite (with at most $k$ parts due to the $K_{k+1}$-freeness).
Lemma 2.3.
Suppose that $F$ is a complete $r$-partite graph with $r\geq 2$, and let $k\geq r$ be an integer. Then for every $n\geq 1$, there exists a complete multipartite graph $G$ on $n$ vertices with at most $k$ parts such that $I(F,G)=I_{k+1}(F,n)$.
Therefore, by Lemmas 2.2 and 2.3, in determining $i(F)$ and $i_{k+1}(F)$, it suffices to restrict our attention to the class of complete multipartite graphs, which, in the limit, reduces to an optimization problem of homogeneous polynomials (see [7, Corollary 6]). Since we also aim to prove perfect stability, we need to consider a slightly more general polynomial (introduced in [22], see (6)) and prove the finiteness of the set of maximizers. For this, we require the following definitions, mostly taken from [22, Section 2].
Define the partite limit space by
$\overline{\mathcal{P}}\coloneqq\Big\{(x_{1},x_{2},\ldots)\in\mathbb{R}^{\mathbb{N}_{+}}\colon\ x_{1}\geq x_{2}\geq\cdots\geq 0\text{ and }\sum_{i\in\mathbb{N}_{+}}x_{i}\leq 1\Big\}.$
For every sequence $\bm{x}=(x_{1},x_{2},\ldots)\in\overline{\mathcal{P}}$, let $x_{0}\coloneqq 1-\sum_{i\in\mathbb{N}_{+}}x_{i}$. Define the support of $\bm{x}$ as
$\mathrm{supp}(\bm{x})\coloneqq\{i\in\mathbb{N}_{+}\colon x_{i}>0\}.$
A sequence in $\overline{\mathcal{P}}$ gives rise to a complete multipartite graph on $n$ vertices via the following definition (the converse is straightforward).
Definition 2.4 ([22, Definition 3]).
Given $\bm{x}=(x_{1},x_{2},\ldots)\in\overline{\mathcal{P}}$ and $n\in\mathbb{N}$, the $n$-vertex realization of $\bm{x}$, denoted $G_{n,\bm{x}}$, is the complete multipartite graph on the vertex set $[n]$ with parts $V_{1},\ldots,V_{m}$ (for some $m$) and a set $V_{0}$ of universal vertices (i.e., the collection of singleton parts of $G_{n,\bm{x}}$), obtained as follows: if $x_{0}=0$, then take a partition $n]=V_{1}\cup\cdots\cup V_{m}$ such that $||V_{i}|-x_{i}n|<1$ for each $i$, and set $V_{0}=\varnothing$; otherwise, for all $i\geq 1$ with $x_{i}n\geq 2$, set $|V_{i}|=\lfloor x_{i}n\rfloor\geq 2$, and let $V_{0}$ contain the remaining vertices in $[n]$. We call $V_{0},\ldots,V_{m}$ the $\overline{\mathcal{P}}$-structure of $G_{n,\bm{x}}$.
For a tuple $(d_{1},\ldots,d_{t})$ of positive integers and a sequence $(x_{1},x_{2},\ldots)\in\mathbb{R}^{\mathbb{N}_{+}}$ of real numbers, let
$S_{d_{1},\ldots,d_{t}}(x_{1},x_{2},\ldots)\coloneqq\sum_{(i_{1},\ldots,i_{t})\in(\mathbb{N}_{+})_{t}}x_{i_{1}}^{d_{1}}\cdots x_{i_{t}}^{d_{t}}.$ (5)
For convenience, we set $S_{\varnothing}(x_{1},x_{2},\ldots)\coloneqq 1$.
For two graphs $F$ and $G$, let $p(F,G)\coloneqq I(F,G)/\binom{v(G)}{v(F)}$ denote the induced density of $F$ in $G$.
Let $a_{1}\geq\cdots\geq a_{r}\geq 1$ be integers, and let $F=K_{a_{1},\ldots,a_{r}}$. Define the limit function
$p_{F}(\bm{x})\coloneqq\lim_{n\to\infty}p(F,G_{n,\bm{x}}).$
It follows from Fact 2.1 that (recall the definition of $\kappa_{F}$ from (2))
$p_{F}(\bm{x})=\kappa_{F}\cdot\sum_{i=0}^{\mathrm{sc}(F)}\binom{\mathrm{sc}(F)}{i}x_{0}^{i}\cdot S_{a_{1},\ldots,a_{r-i}}(\bm{x}).$ (6)
In particular, if $\mathrm{sc}(F)=0$ or $x_{0}=0$, this expression simplifies to
$p_{F}(\bm{x})=\kappa_{F}\cdot S_{a_{1},\ldots,a_{r}}(\bm{x}).$ (7)
Intuitively, Lemma 2.2 and Fact 2.1 together indicate that determining $i(F)$ for a complete multipartite graph amounts to finding the maximum of $p_{F}(\bm{x})$ over $\overline{\mathcal{P}}$. A rigorous proof is given in [22, Section 2].
Theorem 2.5 ([22]).
Suppose that $F$ is a complete multipartite graph. Then
$i(F)=\max_{\bm{x}\in\overline{\mathcal{P}}}~p_{F}(\bm{x}).$
For perfect stability, it is also necessary to understand the property of the maximizers of $p_{F}(\bm{x})$ over $\overline{\mathcal{P}}$. Thus, for a complete multipartite graph $F$, we define
$\mathrm{OPT}(F)\coloneqq\left\{\bm{x}\in\overline{\mathcal{P}}\colon p_{F}(\bm{x})=i(F)\right\}.$
Note that for every $\bm{x}\in\mathrm{OPT}(F)$, the graph $G_{n,\bm{x}}$ is asymptotically extremal for the inducibility problem $I(F,n)$.
Similarly, it follows relatively straightforwardly from Lemma 2.3 and standard results from the theory of graphons (see [32, Section 1] for further details) that, for a complete multipartite graph $F=K_{a_{1},\ldots,a_{r}}$, we have
$i_{k+1}(F)=\max_{\bm{x}\in\mathbb{S}^{k-1}}p_{F}(\bm{x})=\kappa_{F}\cdot\max_{\bm{x}\in\mathbb{S}^{k-1}}S_{a_{1},\ldots,a_{r}}(\bm{x}),$ (8)
where
$\mathbb{S}^{k-1}\coloneqq\left\{(x_{1},\ldots,x_{k},0,\ldots)\in\mathbb{R}^{\mathbb{N}_{+}}\colon\text{$x_{i}\geq 0$ for $i\in[k]$ and $x_{1}+\cdots+x_{k}=1$}\right\}.$
Observe that $\mathbb{S}^{k-1}\subseteq\overline{\mathcal{P}}$ and the union $\bigcup_{k\geq 1}\mathbb{S}^{k-1}$ is the collection of sequences $\bm{x}\in\overline{\mathcal{P}}$ with $x_{0}=0$.
Define
$\mathrm{OPT}_{k+1}(F)\coloneqq\left\{\bm{x}\in\mathbb{S}^{k-1}\colon p_{F}(\bm{x})=i_{k+1}(F)\right\}.$
For convenience, we set $\mathbb{S}^{\infty}=\bigcup_{k\geq 1}\mathbb{S}^{k-1}$, $\mathrm{OPT}_{\infty}(F)=\mathrm{OPT}(F)$, and $i_{\infty}(F)=i(F)$.
For every $k\in\mathbb{N}_{+}$, define
$\bm{k}\coloneqq\big(\underbrace{1/k,\,\ldots,\,1/k}_{k},\,0,\,\ldots\big).$
We call $\bm{x}=(x_{1},x_{2},\ldots)\in\overline{\mathcal{P}}$ balanced if $x_{i}=x_{j}$ for all $i,j\in\mathrm{supp}(\bm{x})$; otherwise it is unbalanced. In particular, if such a balanced $\bm{x}$ belongs to $\mathbb{S}^{k-1}$, then necessarily $\bm{x}=\bm{t}$ for some integer $t\leq k$.
3 Asymptotic results and uniqueness of the maximizer
In this section, we prove the following theorem, which constitutes the most crucial part of the proofs in this work. In particular, it implies Theorem 1.3 and the asymptotic part of Theorem 1.1.
Theorem 3.1.
Let $\ell>r\geq 2$ be integers, and let $m=m_{r,\ell}$. Suppose that $F$ is an almost balanced complete $r$-partite graph on $\ell$ vertices. Then
$\mathrm{OPT}_{k+1}(F)=\begin{cases}\{\bm{k}\},&\quad\text{if}\quad k\in[r,~m-1],\\[1.99997pt] \{\bm{m}\},&\quad\text{if}\quad k\in[m,~\infty].\end{cases}$
In particular,
$i_{k+1}(F)=\begin{cases}\kappa_{F}\cdot\frac{(k-1)_{r-1}}{{k}^{\ell-1}},&\quad\text{if}\quad k\in[r,~m-1],\\[3.00003pt] \kappa_{F}\cdot\frac{(m-1)_{r-1}}{m^{\ell-1}},&\quad\text{if}\quad k\in[m,~\infty].\end{cases}$
In the next subsection, we establish some preliminary inequalities. We then present the proof of Theorem 3.1 in two cases according to the size of $\ell$: $\ell\leq 2r-1$ and $\ell\geq 2r$.
We remark that the proof of the case $\ell\leq 2r-1$ in Theorem 3.1 is similar to that of [32, Theorem 1.6] (which deals with the case $\ell\leq 3r+1$), with one difference: for $k=\infty$, Yuster proved the uniqueness of the maximizer of (7), whereas we prove the uniqueness of the maximizer of (6) (which is slightly stronger but necessary for perfect stability). In contrast, our proof for the case $\ell\geq 2r$ differs substantially from Yuster’s. It relies on a quite nontrivial inequality (Proposition 3.7), whereas Yuster’s argument is based on properties of low-degree polynomials.
3.1 Basic properties of $f$ in (3) and the integer $m_{r,\ell}$
Let $I\subseteq\mathbb{R}$, and let $h\colon I\to\mathbb{R}$ be a function on $I$. We say that $h$ is unimodal if there exists a point $x_{0}\in I$ such that
  • $h$ is non-decreasing on $I\cap(-\infty,x_{0}]$, and
  • $h$ is non-increasing on $I\cap(x_{0},\infty)$.
Lemma 3.2.
Let $\ell>r\geq 2$ be integers. The discrete function $f\colon[r,\infty)\to\mathbb{R}$, defined as in (3), is unimodal and attains its maximum at a unique integer $m_{r,\ell}\in[r,\infty)$.
Proof.
Consider the real extension $\tilde{f}\colon[r,\infty)\to\mathbb{R}$ of $f$ defined by
$\tilde{f}(x)=\frac{(x-1)\cdots(x-r+1)}{x^{\ell-1}},\quad\text{for every real number $x\in[r,\infty)$}.$
Straightforward calculations show that the derivative of $\tilde{f}(x)$ is
$\tilde{f}^{\prime}(x)=\frac{\tilde{f}(x)}{x}\left(\frac{x}{x-1}+\cdots+\frac{x}{x-r+1}-\ell+1\right).$
Define
$g(x)\coloneqq\sum_{i=1}^{r-1}\frac{x}{x-i}-\ell+1,\quad\text{for every real number $x\in[r,\infty)$}.$ (9)
Observe that $g(x)$ is strictly decreasing on $r,\infty)$ and satisfies $\lim_{x\to\infty}g(x)=r-\ell<0$. Thus, the function $\tilde{f}$ is not monotone increasing and is unimodal. It follows that the discrete function $f$ can have at most two maximizers on the set of integers $\{r,r+1,\ldots\}$, and if there are two, they must be consecutive integers.
Suppose, for contradiction, that $f$ attains its maximum at two consecutive integers $k_{0}$ and $k_{0}+1$. Note that, for every $k\in\{r,r+1,\ldots\}$, we have $f(k)<1$, in particular, $f(k)$ is not an integer. Moreover, when written in lowest terms, the denominator of $f(k_{0})$ divides $k_{0}^{\ell-1}$, while the denominator of $f(k_{0}+1)$ divides $(k_{0}+1)^{\ell-1}$. Since $k_{0}$ and $k_{0}+1$ are coprime, $k_{0}^{\ell-1}$ and $(k_{0}+1)^{\ell-1}$ are also coprime, and consequently the denominators of $f(k_{0})$ and $f(k_{0}+1)$ are coprime (and both greater than $1$, since both $f(k_{0})$ and $f(k_{0}+1)$ are not integers). Hence, $f(k_{0})\neq f(k_{0}+1)$, a contradiction. Therefore, the maximizer of $f$ on $\{r,r+1,\ldots\}$ is unique. ∎
It follows from Lemma 3.2 that $m_{r,\ell}=r$ if and only if $f(r)>f(r+1)$, which is equivalent to
$\ell>\frac{\ln(r+1)}{\ln(1+1/r)}=(1+o(1))r\ln r.$
This is consistent with the result of Brown–Sidorenko [7, Theorem 9]. Refined bounds on $m_{r,\ell}$ will be provided in Lemma 4.2.
3.2 Proof of Theorem 3.1 for $\ell<2r$
In this subsection we present the proof of Theorem 3.1 for $\ell<2r$.
Proof of Theorem 3.1 for $\ell<2r$.
Let $F=K_{a_{1},\ldots,a_{r}}$ be an almost balanced complete $r$-partite graph on $\ell\in[r+1,~2r-1]$ vertices (i.e. $\ell=a_{1}+\cdots+a_{r}$), and assume that $a_{1}\geq\cdots\geq a_{r}$. Let $m=m_{r,\ell}$ be the integer give by Lemma 3.2. By the Pigeonhole Principle, we have $a_{r}=1$. By the definition of almost balanced (see (1)), we have $\binom{a_{1}-1}{2}<1$, which means that $a_{1}\leq 2$. Thus we have
$a_{1}=\cdots=a_{\ell-r}=2\quad\text{and}\quad a_{\ell-r+1}=\cdots=a_{r}=1.$
Case 1: $k\in[m,\infty]$.
Since $\mathbb{S}^{k-1}\subseteq\overline{\mathcal{P}}$, it suffices to prove that $\mathrm{OPT}(F)=\{\bm{m}\}$ for this case.
Recall from (5) and (6) that for every $\bm{x}=(x_{1},x_{2},\ldots)\in\overline{\mathcal{P}}$,
$p_{F}(\bm{x})=\kappa_{F}\cdot\sum_{j=0}^{2r-\ell}\tbinom{2r-\ell}{j}x_{0}^{j}\,\sum_{(i_{1},\ldots,i_{r-j})\in(\mathbb{N}_{+})_{r-j}}x_{i_{1}}^{a_{1}}\cdots x_{i_{r-j}}^{a_{r-j}}.$ (10)
Claim 3.3.
The following statements hold.
  1. (i)
    If $\ell\geq r+2$, then every maximizer $\bm{x}\in\mathrm{OPT}(F)$ is balanced.
  2. (ii)
    If $\ell=r+1$, then for every unbalanced maximizer $\bm{x}\in\mathrm{OPT}(F)$, there exists a balanced maximizer $\bm{y}\in\mathrm{OPT}(F)$ such that
    $y_{0}=x_{0}\quad\text{and}\quad|\mathrm{supp}(\bm{y})|<|\mathrm{supp}(\bm{x})|.$
Proof.
Suppose to the contrary that this claim is not true (for both items). Let $\bm{x}=(x_{1},x_{2},\ldots)\in\mathrm{OPT}(F)$ be a counterexample with minimum size of support. Then there exists some distinct pair $i,j\in\mathrm{supp}(\bm{x})$ such that $x_{i}\neq x_{j}$.
Fix all $x_{k}$ where $k\in\mathbb{N}\setminus\{i,j\}$. Note that since $\sum_{k=1}^{\infty}x_{k}=1$, the sum $\beta\coloneqq x_{i}+x_{j}$ is also fixed. From (10), there exist constants $K_{1},\ldots,K_{5}$ independent of $x_{i}$ and $x_{j}$ such that
$p_{F}(\bm{x})$ $=(x_{i}+x_{j})K_{1}+({x_{i}}^{2}+{x_{j}}^{2})K_{2}+x_{i}x_{j}K_{3}+x_{i}x_{j}(x_{i}+x_{j})K_{4}+(x_{i}x_{j})^{2}K_{5}$
$=(x_{i}+x_{j})K_{1}+\left((x_{i}+x_{j})^{2}-2x_{i}x_{j}\right)K_{2}+x_{i}x_{j}K_{3}+x_{i}x_{j}(x_{i}+x_{j})K_{4}+(x_{i}x_{j})^{2}K_{5}.$ (11)
Since $a_{1}=2$ and $a_{r}=1$, we have $\min\{K_{1},K_{2},K_{4}\}>0$. Moreover, $K_{3}>0$ if $a_{r-1}=1$ (equivalently, $\ell\leq 2r-2$), and $K_{5}>0$ if $a_{2}=2$ (equivalently, $\ell\geq r+2$).
Let $\alpha\coloneqq x_{i}x_{j}$. Then $p_{F}(\bm{x})$ can be rewritten as
$\psi(\alpha)\coloneqq p_{F}(\bm{x})$ $=\beta K_{1}+(\beta^{2}-2\alpha)K_{2}+\alpha K_{3}+\alpha\beta K_{4}+\alpha^{2}K_{5}$
$=K_{5}\alpha^{2}+(\beta K_{4}+K_{3}-2K_{2})\alpha+(\beta K_{1}+\beta^{2}K_{2}).$
Now observe that, while $\beta$ is fixed, the product $\alpha=x_{i}x_{j}$ can vary continuously from 0 to $\beta^{2}/4$ by adjusting $x_{i}$ and $x_{j}$ under the constraint $x_{i}+x_{j}=\beta$. Thus we may view $\psi(\alpha)$ as a quadratic polynomial in $\alpha$.
Let $\bm{y}$ and $\bm{z}$ be the sequences in $\overline{\mathcal{P}}$ obtained from $\bm{x}$ by replacing $(x_{i},x_{j})$ with $(x_{i}+x_{j},0)$ and $\left(\frac{x_{i}+x_{j}}{2},\frac{x_{i}+x_{j}}{2}\right)$, respectively, and then reordering the entries accordingly. Note that $y_{0}=z_{0}=x_{0}$.
Suppose that $\ell\geq r+2$. Then $K_{5}>0$. So $\psi(\alpha)$ is a quadratic polynomial in $\alpha$, and hence
$p_{F}(\bm{x})=\psi(\alpha)<\max\left\{\psi(0),~\psi\left(\tfrac{(x_{i}+x_{j})^{2}}{4}\right)\right\}=\max\left\{p_{F}(\bm{y}),~p_{F}(\bm{z})\right\}.$
However, this is a contradiction to the maximality of $p_{F}(\bm{x})$. This proves (i).
Now suppose that $\ell=r+1$. Then $K_{5}=0$. Hence, $\psi(\alpha)$ is linear in $\alpha$ with coefficient $\beta K_{4}+K_{3}-2K_{2}$. If $\beta K_{4}+K_{3}-2K_{2}\neq 0$, then similarly, we have
$p_{F}(\bm{x})=\psi(\alpha)<\max\left\{\psi(0),~\psi\left(\tfrac{(x_{i}+x_{j})^{2}}{4}\right)\right\}=\max\left\{p_{F}(\bm{y}),~p_{F}(\bm{z})\right\},$
again a contradiction. So it must be the case that $\beta K_{4}+K_{3}-2K_{2}=0$. It follows that
$p_{F}(\bm{x})=\psi(\alpha)=\psi(0)=p_{F}(\bm{y}).$
This means that $\bm{y}\in\mathrm{OPT}(F)$ as well. Since $\bm{x}$ does not satisfy Claim 3.3 (ii), the maximizer $\bm{y}$ must be unbalanced. Since $|\mathrm{supp}(\bm{y})|<|\mathrm{supp}(\bm{x})|$, it follows from the minimality of $\bm{x}$ that $\bm{y}$ satisfies Claim 3.3 (ii). That is, there exists a balanced maximizer $\bm{w}=(w_{1},w_{2},\ldots)\in\mathrm{OPT}(F)$ with $w_{0}=y_{0}=x_{0}$ and $|\mathrm{supp}(\bm{w})|<|\mathrm{supp}(\bm{y})|<|\mathrm{supp}(\bm{x})|$. However, $\bm{w}$ witnesses that $\bm{x}$ itself satisfies the assertion of Claim 3.3 (ii), contradicting the assumption that $\bm{x}$ is a counterexample. This completes the proof of Claim 3.3. ∎
Now we establish the following inequality (which is slightly stronger than what we need) for sequences $\bm{x}\in\overline{\mathcal{P}}$ with $x_{0}=0$. Recall that for every $t\in\mathbb{N}_{+}$, $\bm{t}=(1/t,\ldots,1/t,0,\ldots)$.
Claim 3.4.
There exists a constant $\varepsilon=\varepsilon_{r,\ell}>0$, such that for all $\bm{x}\in\overline{\mathcal{P}}$ with $x_{0}=0$, we have
$\Phi(\bm{x})\coloneqq p_{F}(\bm{x})+\varepsilon\bigg(\sum_{(i,j)\in(\mathbb{N}_{+})_{2}}x_{i}x_{j}-\tfrac{m-1}{m}\bigg)^{2}+\varepsilon\bigg(\sum_{(i,j,k)\in(\mathbb{N}_{+})_{3}}x_{i}x_{j}x_{k}-\tfrac{(m-1)(m-2)}{m^{2}}\bigg)^{2}\leq\Phi(\bm{m}),$
and the equality holds if and only if $\bm{x}=\bm{m}$, where $m=m_{r,\ell}$ is the constant given by Lemma 3.2.
Proof.
Recall that for any $t\in\mathbb{N}_{+}$, $p_{F}(\bm{t})=\kappa_{F}\cdot\frac{(t-1)_{r-1}}{k^{\ell-1}}=\kappa_{F}\cdot f(t)$, where $f(t)$ is defined in (3). In what follows, we assume f(r1)=0f(r-1)=0. Define
$\varepsilon\coloneqq\frac{\kappa_{F}}{3}\cdot\min\!\big\{f(m)-f(m-1),\,f(m)-f(m+1)\big\}.$
It follows from Lemma 3.2 that $\varepsilon>0$.
Since $f(x)$ is unimodal (by Lemma 3.2), for every integer $t\neq m$, we have
$\Phi(\bm{t})$ $=\kappa_{F}\cdot f(t)+\varepsilon\left(\tfrac{t-1}{t}-\tfrac{m-1}{m}\right)^{2}+\varepsilon\left(\tfrac{(t-1)(t-2)}{t^{2}}-\tfrac{(m-1)(m-2)}{m^{2}}\right)^{2}$
$\leq\kappa_{F}\cdot\max\big\{f(m-1),~f(m+1)\big\}+\varepsilon+\varepsilon$
$=\kappa_{F}\left(\max\big\{f(m-1),~f(m+1)\big\}+\tfrac{2\kappa_{F}}{3}\cdot\min\big\{f(m)-f(m-1),~f(m)-f(m+1)\big\}\right)$
$<\kappa_{F}\cdot f(m)=\Phi(\bm{m}).$ (12)
Suppose to the contrary that this claim fails. Let $\bm{x}=(x_{1},x_{2},\ldots)\in\overline{\mathcal{P}}\setminus\{\bm{m}\}$ be a counterexample with $x_{0}=0$, i.e., $\Phi(\bm{x})\geq\Phi(\bm{m})$ but $\bm{x}\neq\bm{m}$. Using (3.2), we conclude that there exist $i,j\in\mathrm{supp}(\bm{x})$ such that $x_{i}\neq x_{j}$. Let $\alpha\coloneqq x_{i}x_{j}$ and $\beta\coloneqq x_{i}+x_{j}$. Similar to the proof of Claim 3.3 (more specifically, (3.2)), fix all $x_{k}$ where $k\in\mathbb{N}\setminus\{i,j\}$. Note that since $\sum_{k=1}^{\infty}x_{k}=1$, the sum $\beta=x_{i}+x_{j}$ is also fixed. So we may view $\Phi(\bm{x})$ as a polynomial in $\alpha$:
$\phi(\alpha)\coloneqq\Phi(\bm{x})=\$$p_{F}(\bm{x})+\varepsilon\left(2\alpha+J_{1}\right)^{2}+\varepsilon\big(6\alpha(1-\beta)+J_{2}\big)^{2}$
$=\$$\left(K_{5}+4\varepsilon+36\varepsilon(1-\beta)^{2}\right)\alpha^{2}+\left(\beta K_{4}+K_{3}-2K_{2}+4\varepsilon J_{1}+12\varepsilon(1-\beta)J_{2}\right)\alpha$
$+\left(\beta K_{1}+\beta^{2}K_{2}+\varepsilon J_{1}^{2}+\varepsilon J_{2}^{2}\right),$
where $J_{1},J_{2}$ are constants independent of $x_{i},x_{j}$, and $K_{1},\ldots,K_{5}$ are the same constants as in (3.2).
Let $\bm{y}$ and $\bm{z}$ be the sequences in $\overline{\mathcal{P}}$ obtained from $\bm{x}$ by replacing $(x_{i},x_{j})$ with $(x_{i}+x_{j},0)$ and $\left(\frac{x_{i}+x_{j}}{2},\frac{x_{i}+x_{j}}{2}\right)$, respectively, and then reordering the entries accordingly. Note that $y_{0}=z_{0}=x_{0}=0$. Since $\phi(\alpha)$ is a quadratic polynomial in $\alpha$ with coefficient $K_{5}+4\varepsilon+36\varepsilon(1-\beta)^{2}>0$, we have
$\Phi(\bm{x})=\phi(\alpha)<\max\left\{\phi(0),~\phi\left(\tfrac{(x_{i}+x_{j})^{2}}{4}\right)\right\}=\max\left\{\Phi(\bm{y}),~\Phi(\bm{z})\right\},$
a contradiction to the maximality of $\Phi(\bm{x})$. This completes the proof of Claim 3.4. ∎
Claim 3.5.
Suppose that $\bm{x}\in\overline{\mathcal{P}}$ satisfies $x_{0}>0$. Then $\bm{x}\not\in\mathrm{OPT}(F)$.
Proof.
Suppose to the contrary that this is not true. Then by Claim 3.3, there exists a counterexample $\bm{x}\in\mathrm{OPT}(F)$ with $x_{0}>0$ and $x_{1}=\cdots=x_{s}=y\coloneqq\frac{1-x_{0}}{s}$ for some $s\geq 1$. For every integer $k\geq 1$, let $\bm{x}^{(k)}$ be the sequence obtained from $\bm{x}$ by appending the length-$k$ sequence $\frac{x_{0}}{k},\ldots,\frac{x_{0}}{k}$, and then reordering the entries. By applying (10) to $p_{F}(\bm{x})$ and (7) together with (5) to $p_{F}(\bm{x}^{(k)})$, we obtain that
$p_{F}(\bm{x})-p_{F}(\bm{x}^{(k)})$ $=\kappa_{F}\cdot\sum_{j=0}^{2r-\ell}\tbinom{2r-\ell}{j}(1-sy)^{j}(s)_{r-j}y^{\ell-j}$
$\qquad-\kappa_{F}\cdot\sum_{j=0}^{2r-\ell}\tbinom{2r-\ell}{j}(k)_{j}\left(\tfrac{1-sy}{k}\right)^{j}(s)_{r-j}y^{\ell-j}$
$\qquad-\kappa_{F}\cdot\sum_{i=1}^{\ell-r}\sum_{j=0}^{2r-\ell}\tbinom{\ell-r}{i}\tbinom{2r-\ell}{j}(k)_{i+j}\left(\tfrac{1-sy}{k}\right)^{2i+j}(s)_{r-i-j}y^{\ell-2i-j}.$
Since $\frac{(k)_{j}}{k^{j}}\to 1$ and the third term goes to 0 as $k\to\infty$ (because $\frac{(k)_{i+j}}{k^{2i+j}}\leq\frac{1}{k}\to 0$), we have
$p_{F}(\bm{x})-p_{F}(\bm{x}^{(k)})\to 0\quad\text{as}\quad k\to\infty.$
Combining this with Claim 3.4, we obtain
$0\leq\Phi(\bm{x}^{(k)})-p_{F}(\bm{x}^{(k)})$ $\leq\Phi(\bm{m})-p_{F}(\bm{x}^{(k)})$
$=p_{F}(\bm{m})-p_{F}(\bm{x}^{(k)})\leq p_{F}(\bm{x})-p_{F}(\bm{x}^{(k)})\to 0\quad\text{as}\quad k\to\infty.$
This implies that both
$\lim_{k\to\infty}\sum_{(i,j)\in(\mathbb{N}_{+})_{2}}\bm{x}^{(k)}_{i}\bm{x}^{(k)}_{j}-\tfrac{m-1}{m}=0,\quad\text{and}\qquad\lim_{k\to\infty}\sum_{(i,j,k)\in(\mathbb{N}_{+})_{3}}\bm{x}^{(k)}_{i}\bm{x}^{(k)}_{j}\bm{x}^{(k)}_{k}-\tfrac{(m-1)(m-2)}{m^{2}}=0.$
Simplifying this, we obtain
$\frac{m-1}{m}$ $=(1-sy)^{2}+2(1-sy)sy+s(s-1)y^{2},$
$\frac{(m-1)(m-2)}{m^{2}}$ $=(1-sy)^{3}+3(1-sy)^{2}sy+3(1-sy)s(s-1)y^{2}+s(s-1)(s-2)y^{3}.$
It follows from the first equality that $y=\frac{1}{\sqrt{ms}}$. Plugging it into the second equality we obtain
$1-\frac{3}{m}+\frac{2s}{(ms)^{3/2}}=\frac{(m-1)(m-2)}{m^{2}},$
which implies $s=m$. However, this means that $x_{0}=1-sy=0$, contradicting the choice of $\bm{x}$. ∎
Suppose that $\bm{x}\in\mathrm{OPT}(F)$. It then follows from Claim 3.5 that $x_{0}=0$, and from Claim 3.4 that $\bm{x}=\bm{m}$. This shows that $\mathrm{OPT}(F)=\{\bm{m}\}$ and thereby completes the proof of Case 1.
Case 2: $k\in[r,m-1]$.
An almost identical argument to the proof of Claim 3.3 yields the following result, so we omit the details here.
Claim 3.6.
The following statements hold.
  1. (i)
    If $\ell\geq r+2$, then every maximizer $\bm{x}\in\mathrm{OPT}_{k+1}(F)$ is balanced.
  2. (ii)
    If $\ell=r+1$, then for every unbalanced maximizer $\bm{x}\in\mathrm{OPT}_{k+1}(F)$, there exists a balanced maximizer $\bm{y}\in\mathrm{OPT}_{k+1}(F)$ such that $|\mathrm{supp}(\bm{y})|<|\mathrm{supp}(\bm{x})|$.
Recall that for every integer $t\geq r$, we have $p_{F}(\bm{t})=\kappa_{F}\cdot f(t)$, with $f(t)$ defined in (3). By Lemma 3.2, the function $f$ is strictly increasing on $[r,k]$, so $p_{F}(\bm{t})<p_{F}(\bm{k})$ for every $t<k$. Hence, $i_{k+1}(F)=p_{F}(\bm{k})$.
If $\ell\geq r+2$, then Claim 3.6 (i) immediately implies that $\mathrm{OPT}_{k+1}(F)={\bm{k}}$.
Now, suppose $\ell=r+1$. Take any maximizer $\bm{x}\in\mathrm{OPT}_{k+1}(F)$. If $\bm{x}\neq\bm{k}$, then by Claim 3.6 (ii) we obtain another balanced maximizer $\bm{y}\in\mathrm{OPT}_{k+1}(F)$ with $|\mathrm{supp}(\bm{y})|<|\mathrm{supp}(\bm{x})|$. Consequently, $p_{F}(\bm{y})=p_{F}(\bm{t})$ for some $t<k$, contradicting $i_{k+1}(F)=p_{F}(\bm{k})>p_{F}(\bm{t})$. Therefore, we must have $\bm{x}=\bm{k}$, and again $\mathrm{OPT}_{k+1}(F)=\{\bm{k}\}$.
This completes the proof of Theorem 3.1 for $\ell\leq 2r-1$. ∎
3.3 Proof of Theorem 3.1 for $\ell\geq 2r$
In this subsection, we present the proof of Theorem 3.1 for $\ell\geq 2r$. We will use the following inequality, which constitutes the most crucial ingredient of our proof and distinguishes our approach from those in [7, 5, 23, 32].
Proposition 3.7.
Let $x$ and $y$ be two positive real numbers. For positive integers $d,d_{1},d_{2}$ define
$\mu_{d}$ $\coloneqq(x+y)^{d}-x^{d}-y^{d},\qquad$ $\omega_{d}$ $\coloneqq x^{d}+y^{d}-2\left(\tfrac{x+y}{2}\right)^{d},$
$\psi_{d_{1},d_{2}}$ $\coloneqq x^{d_{1}}y^{d_{2}}+x^{d_{2}}y^{d_{1}},\qquad$ $\phi_{d_{1},d_{2}}$ $\coloneqq 2\left(\tfrac{x+y}{2}\right)^{d_{1}+d_{2}}-x^{d_{1}}y^{d_{2}}-x^{d_{2}}y^{d_{1}}.$
Suppose that $x\neq y$, and let $a,b,s,t$ be positive integers satisfying $b\geq t\geq s\geq a>\binom{b-a}{2}$. Then
$\frac{\omega_{t}}{\mu_{t}}\leq\frac{\omega_{a}}{\mu_{a}}<\frac{\phi_{s,t}}{\psi_{s,t}}.$ (13)
Equivalently,
$\mu_{t}\omega_{a}-\mu_{a}\omega_{t}\geq 0,\quad\text{and}\qquad\mu_{a}\phi_{s,t}-\omega_{a}\psi_{s,t}>0.$ (14)
Proof.
First we prove that $\frac{\omega_{t}}{\mu_{t}}\leq\frac{\omega_{a}}{\mu_{a}}$, which follows from the following claim.
Claim 3.8.
We have $\frac{\omega_{k+1}}{\mu_{k+1}}\leq\frac{\omega_{k}}{\mu_{k}}$ for every integer $k\geq a$.
Proof.
By definition, this is equivalent to showing
$\left(x^{k+1}+y^{k+1}-2\left(\tfrac{x+y}{2}\right)^{k+1}\right)\left((x+y)^{k}-x^{k}-y^{k}\right)\\ \leq\left(x^{k}+y^{k}-2\left(\tfrac{x+y}{2}\right)^{k}\right)\left((x+y)^{k+1}-x^{k+1}-y^{k+1}\right).$
Simplifying this reduces to showing that:
$(x+y)^{k+1}\leq x^{k+1}+y^{k+1}+(2^{k}-1)(x^{k}y+xy^{k}).$ (15)
Since for every $i\in[k]$ it holds that
$(x^{k}y+xy^{k})-(x^{i}y^{k+1-i}+x^{k+1-i}y^{i})=xy(x^{i-1}-y^{i-1})(x^{k-i}-y^{k-i})\geq 0,$
it follows that
$\sum_{i=1}^{k}\tbinom{k+1}{i}(x^{k}y+xy^{k})$ $\geq\sum_{i=1}^{k}\tbinom{k+1}{i}x^{i}y^{k+1-i}+\sum_{i=1}^{k}\tbinom{k+1}{i}x^{k+1-i}y^{i},$
which implies that
$\left(2^{k+1}-2\right)(x^{k}y+xy^{k})\geq 2\left((x+y)^{k+1}-\left(x^{k+1}+y^{k+1}\right)\right)$
After dividing both sides by $2$ and rearranging, we obtain (15). ∎
Next, we prove that $\frac{\omega_{a}}{\mu_{a}}<\frac{\phi_{s,t}}{\psi_{s,t}}$ for $b\geq t\geq s\geq a\geq 2$. By definition, this inequality is equivalent to
$\big((x+y)^{a}-x^{a}-y^{a}\big)\left(2\left(\tfrac{x+y}{2}\right)^{s+t}-x^{s}y^{t}-x^{t}y^{s}\right)>\left(x^{a}+y^{a}-2\left(\tfrac{x+y}{2}\right)^{a}\right)\big(x^{s}y^{t}+x^{t}y^{s}\big).$
Simplifying this reduces to showing that:
$\left(\tfrac{x+y}{2}\right)^{s+t-a}\big((x+y)^{a}-x^{a}-y^{a}\big)>\left(2^{a-1}-1\right)\left(x^{s}y^{t}+x^{t}y^{s}\right).$ (16)
We first prove the following auxiliary inequalities:
Claim 3.9.
The following inequalities hold:
$2^{a-2}\big((x+y)^{a}-x^{a}-y^{a}\big)\geq(2^{a}-2)xy(x+y)^{a-2},$ (17)
and
$2\left(\tfrac{x+y}{2}\right)^{s+t-2}>x^{s-1}y^{t-1}+x^{t-1}y^{s-1}.$ (18)
Proof.
First, we prove (17). It is straightforward to verify that (17) holds with equality when $a\in\{2,3\}$, so we may assume that $a\geq 4$. For $i\in[a-1]$, let
$M_{i}\coloneqq 2^{a-2}\tbinom{a}{i}-(2^{a}-2)\tbinom{a-2}{i-1}$,
noting that $M_{i}=M_{a-i}$. Let
$f(i)\coloneqq\tfrac{2^{a-2}\tbinom{a}{i}}{(2^{a}-2)\tbinom{a-2}{i-1}}=\tfrac{2^{a-2}a(a-1)}{(2^{a}-2)i(a-i)}.$
Observe that $f(i)$ is decreasing in $i$ for $i\in(0,a/2)$ and increasing in $i$ for $i\in(a/2,a)$. Also
$f(1)=f(a-1)=\tfrac{2^{a-2}a}{2^{a}-2}>1\quad\text{and}\quad f(a/2)=\tfrac{2^{a-2}(a-1)}{(2^{a}-2)a}<1.$
So there exists $i_{0}\in[2,a/2]$ such that
[…]
For $i,j\in[1,a/2]$ with $i<j$, we have
$\left(x^{i}y^{a-i}+x^{a-i}y^{i}\right)-\left(x^{j}y^{a-j}+x^{a-j}y^{j}\right)=x^{i}y^{i}\left(x^{j-i}-y^{j-i}\right)\left(x^{a-i-j}-y^{a-i-j}\right)\geq 0.$
Define $\gamma_{i}\coloneqq x^{i}y^{a-i}$ and $\beta_{i}\coloneqq(\gamma_{i}+\gamma_{a-i})/2$ for $i\in[1,a/2]$. The inequality above shows that for $i<j\leq a/2$, we have $\gamma_{i}+\gamma_{a-i}\geq\gamma_{j}+\gamma_{a-j}$ and consequently $\beta_{i}\geq\beta_{j}$. Also note that
$\sum_{i=1}^{a-1}M_{i}=2^{a-2}\sum_{i=1}^{a-1}\tbinom{a}{i}-\left(2^{a}-2\right)\sum_{i=1}^{a-1}\tbinom{a-2}{i-1}=2^{a-2}\left(2^{a}-2\right)-\left(2^{a}-2\right)2^{a-2}=0,$
Then, using the Binomial Theorem and the symmetry $M_{i}=M_{a-i}$ and $\beta_{i}=\beta_{a-i}$, we obtain that
$2^{a-2}\big((x+y)^{a}-x^{a}-y^{a}\big)-(2^{a}-2)xy(x+y)^{a-2}$
$=$ $\sum_{i=1}^{a-1}M_{i}\gamma_{i}=\frac{1}{2}\sum_{i=1}^{a-1}\left(M_{i}\gamma_{i}+M_{a-i}\gamma_{a-i}\right)=\sum_{i=1}^{a-1}M_{i}\beta_{i}=\sum_{i=1}^{i_{0}-1}M_{i}\beta_{i}+\sum_{i=i_{0}}^{a-i_{0}}M_{i}\beta_{i}+\sum_{i=a-i_{0}+1}^{a-1}M_{i}\beta_{i}$
$\geq$ $\left(\sum_{i=1}^{i_{0}-1}M_{i}\right)\min_{1\leq i\leq i_{0}-1}\beta_{i}+\left(\sum_{i=i_{0}}^{a-i_{0}}M_{i}\right)\max_{i_{0}\leq i\leq a-i_{0}}\beta_{i}+\left(\sum_{i=a-i_{0}+1}^{a-1}M_{i}\right)\min_{a-i_{0}+1\leq i\leq a-1}\beta_{i}$
$=$ $\left(\sum_{i=1}^{i_{0}-1}M_{i}\right)\beta_{i_{0}-1}+\left(\sum_{i=i_{0}}^{a-i_{0}}M_{i}\right)\beta_{i_{0}}+\left(\sum_{i=a-i_{0}+1}^{a-1}M_{i}\right)\beta_{a-i_{0}+1}=2\left(\sum_{i=1}^{i_{0}-1}M_{i}\right)\left(\beta_{i_{0}-1}-\beta_{i_{0}}\right)\geq 0,$
where the last equation uses the symmetry and the fact that $\sum_{i=1}^{a-1}M_{i}=0$. This proves (17).
It remains to prove (18). The argument is adapted from the proof of [7, Theorem 3], with some modifications. Define, for every $z\in[0,1]$,
$g(z)\coloneqq z^{t-1}(1-z)^{s-1}+z^{s-1}(1-z)^{t-1}$
We first show that $g$ attains its maximum uniquely at $z=1/2$.
Letting $x\coloneqq(1-z)/z$. Straightforward calculations show that
$g^{\prime}(z)=z^{t-1}(1-z)^{s-2}\Bigl((s-1)x^{t-s+1}-(t-1)x^{t-s}+(t-1)x-s+1\Bigr),$
Define
$h(x)\coloneqq(s-1)x^{t-s+1}-(t-1)x^{t-s}+(t-1)x-s+1$
Let $y\coloneqq x-1$. Then for $x\geq 1$, we have
$h(x)$ $=(s-1)(1+y)^{t-s+1}-(t-1)(1+y)^{t-s}+(t-1)(1+y)-s+1$
$=\Bigl((s-1)(t-s+1)-(t-1)(t-s)+(t-1)\Bigr)y+\sum_{i=2}^{t-s+1}\bigl((s-1)\tbinom{t-s+1}{i}-(t-1)\tbinom{t-s}{i}\bigr)y^{i}$
$=2\Bigl(s-1-\tbinom{t-s}{2}\Bigr)y+\sum_{i=2}^{t-s+1}\tbinom{t-s}{i-1}\frac{1}{i}\Bigl((s-1)(t-s+1)-(t-1)(t-s-i+1)\Bigr)y^{i}$
$\geq 2\Bigl(s-1-\tbinom{t-s}{2}\Bigr)y+\sum_{i=2}^{t-s+1}\tbinom{t-s}{i-1}\frac{1}{i}\Bigl((s-1)(t-s+1)-(t-1)(t-s-1)\Bigr)y^{i}$
$=2\Bigl(s-1-\tbinom{t-s}{2}\Bigr)y+\sum_{i=2}^{t-s+1}\tbinom{t-s}{i-1}\frac{2}{i}\Bigl(s-1-\tbinom{t-s}{2}\Bigr)y^{i}\geq 0,$
[…]
It follows that $g^{\prime}(z)\geq 0$ for $z\in(0,1/2)$, and hence, $g$ is nondecreasing on $[0,1/2]$. Since $g$ is symmetric around $z=1/2$ and is a polynomial which is not linear (since $s,t\geq 2$), it attains its maximum on the interval $[0,1]$ uniquely at $z=1/2$. Therefore, (18) holds. ∎
Combining (17) and (18), we obtain
$\left(\tfrac{x+y}{2}\right)^{s+t-a}\big((x+y)^{a}-x^{a}-y^{a}\big)$ $\geq\left(\tfrac{x+y}{2}\right)^{s+t-a}\left(2^{a}-2\right)xy\left(\tfrac{x+y}{2}\right)^{a-2}$
$>\tfrac{2^{a}-2}{2}xy\left(x^{s-1}y^{t-1}+x^{t-1}y^{s-1}\right)=\left(2^{a-1}-1\right)\left(x^{s}y^{t}+x^{t}y^{s}\right),$
which proves (16). This completes the proof of Proposition 3.7. ∎
We are now ready to prove Theorem 3.1 for $\ell\geq 2r$.
Proof of Theorem 3.1 for $\ell\geq 2r$.
Let $F=K_{a_{1},\ldots,a_{r}}$ be an almost balanced complete $r$-partite graph on $\ell\geq 2r$ vertices, that is, $\ell=a_{1}+\cdots+a_{r}$. Let $m=m_{r,\ell}$ be the integer give by Lemma 3.2. Assume that $a_{1}\geq\cdots\geq a_{r}$. Since $F$ is almost balanced, we have $a_{r}\geq 2$, otherwise by (1), we would have $a_{1}\leq 2$, which implies that $\ell\leq 2r-1$, a contradiction. Fix $k\in[r,\infty]$.
Claim 3.10.
For every $\bm{x}\in\mathrm{OPT}_{k}(F)$, we have $x_{0}=0$.
Proof.
It follows from the definition of $\mathbb{S}^{k-1}$ that $x_{0}=0$ if $k\neq\infty$. Thus we may assume that $k=\infty$. Suppose to the contrary that there exists $\bm{x}=(x_{1},x_{2},\ldots)\in\mathrm{OPT}(F)$ with $x_{0}>0$. Since in every realization $G_{n,\bm{{x}}}$ (Definition 2.4), there is no induced copy of $F$ containing vertices from $V_{0}$ (because $a_{1}\geq\cdots\geq a_{r}\geq 2$), we have
$p_{F}(\bm{x})=\kappa_{F}\cdot\sum_{(i_{1},\ldots,i_{r})\in(\mathbb{N}_{+})_{r}}x_{i_{1}}^{a_{1}}\cdots x_{i_{r}}^{a_{r}}.$
Let $\tilde{\bm{x}}$ be the sequence obtained from $\bm{x}$ by replacing $x_{1}$ with $x_{0}+x_{1}$. Note that $\tilde{\bm{x}}\in\overline{\mathcal{P}}$ and $p_{F}(\tilde{\bm{x}})-p_{F}(\bm{x})>0$, a contradiction to the optimality of $\bm{x}$. ∎
Fix an arbitrary optimal sequence $\bm{x}\in\mathrm{OPT}_{k}(F)$. By Claim 3.10, we have $x_{0}=0$. We aim to show that all nonzero entries of $\bm{x}$ are equal. Suppose to the contrary that there exist $i,j\in\mathrm{supp}(\bm{x})$ with $x_{i}\neq x_{j}$.
Let $\bm{y}$ and $\bm{z}$ be the sequence in $\overline{\mathcal{P}}$ obtained from $\bm{x}$ by replacing $(x_{i},x_{j})$ with $(x_{i}+x_{j},0)$ and $\left(\frac{x_{i}+x_{j}}{2},\frac{x_{i}+x_{j}}{2}\right)$, respectively, and then reordering the entries accordingly. Note that $y_{0}=z_{0}=x_{0}=0$ and $\{\bm{y},\bm{z}\}\subseteq\mathbb{S}^{k-1}$. In addition, define $\tilde{\bm{x}}$ by replacing both $x_{i}$ and $x_{j}$ with 0 and then reordering the entries.
Let $\bm{a}\coloneqq(a_{1},\ldots,a_{r})$. Recall the definition of $S_{\bm{a}}(\bm{x})$ from (5). We shall prove the following key inequality:
Claim 3.11.
We have
$S_{\bm{a}}(\bm{x})<\max\big\{S_{\bm{a}}(\bm{y}),~S_{\bm{a}}(\bm{z})\}.$ (19)
Consequently, $p_{F}(\bm{x})<\max\big\{p_{F}(\bm{y}),~p_{F}(\bm{z})\}$.
Proof.
For $p,q\in[r]$, let $\bm{a}^{p}$ be the (r1)(r-1)-tuple obtained from $\bm{a}$ by removing the element $a_{p}$; let $\bm{a}^{p,q}$ be the (r2)(r-2)-tuple obtained from $\bm{a}$ by removing the elements $a_{p}$ and $a_{q}$; and let
[…]
It is clear that $|\mathrm{supp}(\bm{x})|\geq r$. Thus, $|\mathrm{supp}(\tilde{\bm{x}})|\geq r-2$, and hence, $B_{p,q}>0$ for all $\{p,q\}\subseteq[r]$.
Define
$\mu_{p}$ $\coloneqq(x_{i}+x_{j})^{a_{p}}-x_{i}^{a_{p}}-x_{j}^{a_{p}},\qquad$ $\omega_{p}$ $\coloneqq x_{i}^{a_{p}}+x_{j}^{a_{p}}-2\left(\tfrac{x_{i}+x_{j}}{2}\right)^{a_{p}},$
$\psi_{p,q}$ $\coloneqq x_{i}^{a_{p}}x_{j}^{a_{q}}+x_{i}^{a_{q}}x_{j}^{a_{p}},\qquad$ $\phi_{p,q}$ $\coloneqq 2\left(\tfrac{x_{i}+x_{j}}{2}\right)^{a_{p}+a_{q}}-x_{i}^{a_{p}}x_{j}^{a_{q}}-x_{i}^{a_{q}}x_{j}^{a_{p}}.$
Suppose to the contrary that this claim fails. Then we have
$0\leq S_{\bm{a}}(\bm{x})-S_{\bm{a}}(\bm{y})$ $=-\sum_{p\in[r]}\mu_{p}A_{p}+\sum_{\{p,q\}\in\binom{[r]}{2}}\psi_{p,q}B_{p,q},$ (20)
$0\leq S_{\bm{a}}(\bm{x})-S_{\bm{a}}(\bm{z})$ $=\sum_{p\in[r]}\omega_{p}A_{p}-\sum_{\{p,q\}\in\binom{[r]}{2}}\phi_{p,q}B_{p,q}.$ (21)
[…]
$\sum_{p\in[r]}(\mu_{p}\omega_{r}-\mu_{r}\omega_{p})A_{p}+\sum_{\{p,q\}\in\binom{[r]}{2}}(\mu_{r}\phi_{p,q}-\omega_{r}\psi_{p,q})B_{p,q}\leq 0.$ (22)
However, it follows from (14) and $B_{p,q}>0$ that
$\sum_{p\in[r]}(\mu_{p}\omega_{r}-\mu_{r}\omega_{p})A_{p}+\sum_{\{p,q\}\in\binom{[r]}{2}}(\mu_{r}\phi_{p,q}-\omega_{r}\psi_{p,q})B_{p,q}>0,$
a contradiction to (22). ∎
It follows from Claim 3.11 and the optimality of $\bm{x}$ that $\bm{x}$ must be balanced. Suppose that $|\mathrm{supp}(\bm{x})|=\hat{k}$ for some integer $\hat{k}\leq k$. Then $\bm{x}=\bm{\hat{k}}=(1/\hat{k},\ldots,1/\hat{k},0,\ldots)$, and hence, $p_{F}(\bm{x})=\kappa_{F}\cdot{(\hat{k})_{r}}/{\hat{k}^{r}}$. It follows from Lemma 3.2 and the maximality of $p_{F}(\bm{x})$ that $\hat{k}=m$ if $k\geq m$ and $\hat{k}=k$ if $k\leq m-1$. This completes the proof of Theorem 3.1 for $\ell\geq 2r$. ∎
4 Perfect stability
4.1 Preparations
In this subsection, we briefly review the sufficient conditions for perfect stability in the framework developed by [22]. Their framework applies to the broader class of “symmetrizable” functions, but here we focus specifically on the inducibility problem for complete multipartite graphs $F$. Very roughly, [22, Theorem 1.1] asserts that if the set of maximizers $\bm{x}$ satisfies certain strictness conditions, then the inducibility problem for $F$ is perfectly stable.
For convenience, we will use the following simplified variant of [22, Theorem 1.1], which is roughly [3, Theorem 8.2] (see also [4, Theorem 5.2]).
Theorem 4.1.
Let $F$ be a complete multipartite graph with $\ell$ vertices. Suppose that $\mathrm{OPT}(F)$ contains a unique sequence $\bm{x}$ with $x_{0}=0$ (i.e. $\sum_{i\in\mathbb{N}_{+}}x_{i}=1$) and $t\coloneqq|\mathrm{supp}(\bm{x})|<\infty$. Suppose that there exist constants $\varepsilon>0$ and $N_{0}$ such that the following holds for the complete multipartite graph $G\coloneqq G_{n,\bm{x}}$ with parts $V_{1},\ldots,V_{t}$ for every $n\geq N_{0}\colon$
  1. (S1)
    For every pair $\{x,y\}\subseteq V(G)$, we have
    $I(F,G)-I(F,G\oplus xy)\geq\varepsilon n^{\ell-2},$
    where $G\oplus xy$ denotes the graph with vertex set $V(G)$ and edge set $E(G)\bigtriangleup\{xy\}$.
  2. (S2)
    For every $A\subseteq[t]$ with $|A|\neq t-1$ and every $A^{*}\subseteq[t]$ with $|A^{*}|=t-1$,
    $I(F,G_{A^{*}})-I(F,G_{A})\geq\varepsilon n^{\ell-1},$
    where for a subset $B\subseteq[t]$, $G_{B}$ denotes the graph obtained from $G$ by adding a new vertex adjacent precisely to $\bigcup_{i\in B}V_{i}$.
Then the inducibility problem for $F$ is perfectly stable.
Next, we derive refined estimates for $m_{r,\ell}$, the integer given by Lemma 3.2.
Lemma 4.2.
Let $\ell>r\geq 2$ be integers and let $m=m_{r,\ell}$ be the unique integer from Lemma 3.2. Then the following inequalities hold:
  1. (i)
    $m>\max\left\{\frac{\ell(r-1)}{2(\ell-r)},~r-1\right\}$.
  2. (ii)
    $m<\min\left\{\frac{\ell(\ell-1)}{2(\ell-r)},~\frac{r}{\alpha}\right\}$, where $\alpha\coloneqq\alpha(\ell/r)\in(1-r^{2}/\ell^{2},1)$ is the unique positive root of the equation $\mathrm{e}^{(\ell/r)\cdot x}(1-x)=1$. In particular, $m<\frac{r\ell^{2}}{\ell^{2}-r^{2}}$.
Proof.
First, we prove (i). Since we trivially have $m\geq r$, it suffices to show that $m>\frac{\ell(r-1)}{2(\ell-r)}$. Note that $\frac{\ell(r-1)}{2(\ell-r)}$ is decreasing in $\ell>r$, and when $\ell=2r$, we have $\frac{\ell(r-1)}{2(\ell-r)}<r$. Thus we may assume that $r<\ell<2r$.
Since $f(x)$ (as defined in (3)) is unimodal, the optimality of $m$ implies that $f(m)>f(m+1)$. Simplifying this inequality yields
$\left(1-\frac{1}{m+1}\right)^{\ell}<1-\frac{r}{m+1}.$
Define $g(x)\coloneqq(1-x)^{\ell}+rx-1$ for $x\in[0,1]$. Then the inequality above implies that $g\left(\tfrac{1}{m+1}\right)<0$.
Claim 4.3.
We have $g\left(\frac{2(\ell-r)}{r\ell+\ell-2r}\right)\geq 0$.
Proof.
It is equivalent to show that
$\left(1-\frac{2(\ell-r)}{r\ell+\ell-2r}\right)^{\ell}\geq 1-\frac{2r(\ell-r)}{r\ell+\ell-2r}.$
Taking logarithms, it simplifies to showing that
$\ell\ln\ell+(\ell-1)\ln(r-1)-(\ell-1)\ln((r+1)\ell-2r)-\ln(2r-\ell)\geq 0.$ (23)
Define
$\vartheta(x)\coloneqq x\ln x+(x-1)\ln(r-1)-(x-1)\ln\big((r+1)x-2r\big)-\ln(2r-x),\quad\text{for}\quad x\in[r,2r).$
Straightforward calculations show that $\vartheta(r)=\vartheta^{\prime}(r)=0$ and
$\vartheta^{\prime\prime}(x)=\frac{4r^{2}(x-r)\bigl((r+2)x-4r\bigr)}{\bigl((r+1)x-2r\bigr)^{2}(2r-x)^{2}x}.$
For $x\in[r,2r)$ and $r\geq 2$, we have $(r+2)x-4r\geq(r+2)r-4r=r(r-2)\geq 0$. Hence $\vartheta^{\prime\prime}(x)\geq 0$. It follows that $\vartheta(x)\geq 0$ for all $x\in[r,2r)$, which implies (23). This completes the proof of Claim 4.3. ∎
Since the second derivative $g^{\prime\prime}(x)=\ell(\ell-1)(1-x)^{\ell-2}\geq 0$, the function $g(x)$ is convex. Note that $g(0)=0$ and $g\left(\frac{1}{m+1}\right)<0$. Combining this with Claim 4.3, we obtain $\frac{1}{1+m}<\frac{2(\ell-r)}{r\ell+\ell-2r}$, which implies that $m>\frac{\ell(r-1)}{2(\ell-r)}$. This proves (i).
It remains to prove (ii). Let $t=\ell/r$. First we show that the real number $\alpha$ defined in the lemma lies in the interval $(1-1/t^{2},1)$. Define $h_{t}(x)\coloneqq\mathrm{e}^{tx}(1-x)-1$. Note that the derivative $h_{t}^{\prime}(x)=\mathrm{e}^{tx}(t-1-tx)$ is positive when x<11/tx<1-1/t and negative when x>11/tx>1-1/t. Thus, $h_{t}(x)$ is unimodal on $(0,\infty)$. Since $h_{t}(0)=0$ and $h_{t}(1)=-1$, the equation $h_{t}(x)=1$ has a unique positive root, which lies in the interval $(0,1)$.
To show that $\alpha>1-1/t^{2}$, it suffices to verify that the following quantity is positive:
$h_{t}(1-1/t^{2})=\frac{\mathrm{e}^{t-\frac{1}{t}}}{t^{2}}-1.$
Let $u(t)\coloneqq\mathrm{e}^{t-\frac{1}{t}}/t^{2}$ for $t>1$. Since the derivative $u^{\prime}(t)=u(t)(1-1/t)^{2}$ is positive, the function $u(t)$ is strictly increasing. Consequently, $h_{t}(1-1/t^{2})=u(t)-1>u(1)-1=0$, as desired.
Next, we prove the upper bound for $m$. Since $\alpha<1$ and straightforward calculations show that $\frac{\ell(\ell-1)}{2(\ell-r)}>r$ for all $\ell>r$, we may assume that $m>r$ (otherwise, we are done).
Since $f(x)$ (as defined in (3)) is unimodal, the optimality of $m$ implies that f(m)>f(m1)f(m)>f(m-1). Simplifying this inequality yields
$\frac{m-r}{m}<\left(1-\frac{1}{m}\right)^{\ell}.$ (24)
Combining it with the inequality
$\left(1-\frac{1}{m}\right)^{\ell}\leq 1-\frac{\ell}{m}+\binom{\ell}{2}\left(\frac{1}{m}\right)^{2},$
we obtain $m<\frac{\ell(\ell-1)}{2(\ell-r)}$.
It remains to prove that $m<r/\alpha$. Define $\theta\coloneqq r/m$. Then it follows from (24) and the inequality $\ln(1+x)\leq x$ for x>1x>-1 that
$1-\theta=\frac{m-r}{m}<\left(1-\frac{1}{m}\right)^{\ell}=\exp\left(\ell\ln\left(1-\frac{1}{m}\right)\right)\leq\exp\left(-\frac{\ell}{m}\right)=\mathrm{e}^{-t\theta}.$
In other words, we have $h_{t}(\theta)<0$. Since $h_{t}(x)$ is unimodal and $h_{t}(0)=0$, we have $\theta>\alpha$, which implies that $m<r/\alpha$. This proves (ii), completing the proof of Lemma 4.2. ∎
4.2 Proof of Theorem 1.5 for $\ell<2r$
In this subsection, we present the proof of Theorem 1.5 for $\ell<2r$.
Proof of Theorem 1.5 for $\ell<2r$.
Let $F=K_{a_{1},\ldots,a_{r}}$ be an almost balanced complete $r$-partite graph on $\ell\in[r+1,~2r-1]$ vertices (i.e. $\ell=a_{1}+\cdots+a_{r}$). Let $m=m_{r,\ell}$ be the integer given by Lemma 3.2. Assume that $a_{1}\geq\cdots\geq a_{r}$. By the Pigeonhole Principle, we have $a_{r}=1$. By the definition of almost balanced (see (1)), we have $\binom{a_{1}-1}{2}<1$, which means that $a_{1}\leq 2$. Thus we have
$a_{1}=\cdots=a_{\ell-r}=2\quad\text{and}\quad a_{\ell-r+1}=\cdots=a_{r}=1.$
It follows from Theorem 3.1 that $\mathrm{OPT}(F)=\{\bm{m}\}$. Thus, by Theorem 4.1, it suffices to verify (S1) and (S2) with some constant $\varepsilon>0$ and all sufficiently large $n$.
Fix a sufficiently large $n$. For simplicity, we will use $o(\cdot)$ to denote lower-order terms and assume that $n=km$ for some integer $k$ (i.e., $n$ is divisible by $m$). It will be clear from the proof that this divisibility assumption has no substantive effect; it is included only to allow us to omit floor and ceiling signs.
Let $G\coloneqq G_{n,\bm{m}}$ be the $n$-vertex realization with parts $V_{1},\ldots,V_{m}$, where $|V_{i}|=k$ for all $i\in[m]$. Let $\varepsilon>0$ be the constant satisfying
$1/\varepsilon=m^{\ell+1}{2^{\ell-r}(\ell-r)!(2r-\ell)!}.$
We first verify (S1) for all distinct $x,y\in V(G)$.
Claim 4.4.
Suppose that $(x,y)\in V_{i}\times V_{j}$ for distinct $i,j\in[m]$. Then
$I(F,G)-I(F,G\oplus xy)\geq\varepsilon n^{\ell-2}.$
Proof.
By symmetry, we may assume that $(x,y)\in V_{1}\times V_{2}$. Since $x$ and $y$ lie in different parts, the pair $\{x,y\}$ is an edge of $G$.
Every induced copy $T$ of $F$ in $G$ that contains the edge $xy$ can be classified into three cases, depending on how its remaining vertices are distributed among the parts $V_{1},\ldots,V_{m}$}.
Case 1: $|V_{1}\cap V(T)|=|V_{2}\cap V(T)|=1$.
If $\ell=2r-1$, the number of the copies of these $T$ is 0, otherwise this number is
$\frac{(m-2)_{r-2}\cdot k^{2r-\ell-2}\cdot\binom{k}{2}^{\ell-r}}{(\ell-r)!(2r-\ell-2)!}=\frac{(m-2)_{r-2}\bigl(k^{\ell-2}+o(k^{\ell-2})\bigr)}{2^{\ell-r}(\ell-r)!(2r-\ell-2)!}.$ (25)
To see this, note that $x$ and $y$ are the unique vertices from $V_{1}$ and $V_{2}$ in $T$. The factor $(m-2)_{r-2}$ indicates the number of ways to choose the r2r-2 other parts of $T$ from the remaining m2m-2 parts (of size $k$). Among the r2r-2 other parts of $T$, there are $2r-\ell-2$ parts of size $1$ and $\ell-r$ parts of size $2$, so the denominator $(\ell-r)!(2r-\ell-2)!$ accounts for overcounting due to permutations of parts of equal size. If $\ell=2r-1$, no part apart from $V_{1},V_{2}$ can have size $1$, so the count is 0.
Using similar estimations, we can obtain the remaining two cases.
Case 2: $|V_{1}\cap V(T)|=1,|V_{2}\cap V(T)|=2$ or $|V_{1}\cap V(T)|=2,|V_{2}\cap V(T)|=1$.
The number of the induced copies of these $T$ is
$2\cdot\frac{(m-2)_{r-2}\cdot(k-1)k^{2r-\ell-1}\cdot\binom{k}{2}^{\ell-r-1}}{(\ell-r-1)!(2r-\ell-1)!}=\frac{(m-2)_{r-2}\bigl(k^{\ell-2}+o(k^{\ell-2})\bigr)}{2^{\ell-r-2}(\ell-r-1)!(2r-\ell-1)!}.$
Case 3: $|V_{1}\cap V(T)|=|V_{2}\cap V(T)|=2$. If $\ell=r+1$, the number of the induced copies of these $T$ is 0, otherwise this number is
$\frac{(m-2)_{r-2}\cdot(k-1)^{2}k^{2r-\ell}\cdot\binom{k}{2}^{\ell-r-2}}{(\ell-r-2)!(2r-\ell)!}=\frac{(m-2)_{r-2}\bigl(k^{\ell-2}+o(k^{\ell-2})\bigr)}{2^{\ell-r-2}(\ell-r-2)!(2r-\ell)!}.$
Summing over all cases (regardless of the relative sizes of $\ell$ and $r$), the number of $T$ containing $xy$ is
$\frac{(m-2)_{r-2}\bigl(k^{\ell-2}+o(k^{\ell-2})\bigr)}{2^{\ell-r}(\ell-r)!(2r-\ell)!}\cdot\bigl((2r-\ell)(2r-\ell-1)+4(\ell-r)(2r-\ell)+4(\ell-r)(\ell-r-1)\bigr)$
$={}$ $\frac{(m-2)_{r-2}(\ell^{2}+2r-3\ell)}{2^{\ell-r}(\ell-r)!(2r-\ell)!}\bigl(k^{\ell-2}+o(k^{\ell-2})\bigr).$
Now we turn to the induced copy $T$ of $F$ in $G\oplus xy$ containing non-edge $xy$, so $x,y$ must constitute a part of $T$. Then the number of the induced copies of these $T$ is
$\frac{(m-2)_{r-1}k^{2r-\ell}\binom{k}{2}^{\ell-r-1}}{(\ell-r-1)!(2r-\ell)!}=\frac{(m-2)_{r-1}\bigl(k^{\ell-2}+o(k^{\ell-2})\bigr)}{2^{\ell-r-1}(\ell-r-1)!(2r-\ell)!}.$
Hence we obtain
$I(F,G)-I(F,G\oplus xy)=$ $\frac{(m-2)_{r-2}(\ell^{2}+2r-3\ell-2(m-r)(\ell-r))}{2^{\ell-r}(\ell-r)!(2r-\ell)!}\bigl(k^{\ell-2}+o(k^{\ell-2})\bigr)$
$=$ $\frac{(m-2)_{r-2}(\ell^{2}+2r-3\ell-2(m-r)(\ell-r))}{2^{\ell-r}m^{\ell-2}(\ell-r)!(2r-\ell)!}\bigl(n^{\ell-2}+o(n^{\ell-2})\bigr).$
By Lemma 4.2 (ii), we have $m<\frac{\ell(\ell-1)}{2(\ell-r)}$. Substituting this upper bound gives
$\ell^{2}+2r-3\ell-2(m-r)(\ell-r)$ $>\ell^{2}+2r-3\ell-2\left(\frac{\ell(\ell-1)}{2(\ell-r)}-r\right)(\ell-r)$
$=\ell^{2}+2r-3\ell-\ell(\ell-1)+2r(\ell-r)$
$=2(r-1)(\ell-r)\geq 2.$
Hence $I(F,G)-I(F,G\oplus xy)>\varepsilon n^{\ell-2}$ for all sufficiently large $n$. This completes the proof of Claim 4.4. ∎
Claim 4.5.
Suppose that $\{x,y\}\subseteq V_{i}$ for some $i\in[m]$. Then
$I(F,G)-I(F,G\oplus xy)\geq\varepsilon n^{\ell-2}.$
Proof.
Assume without loss of generality that $x,y\in V_{1}$ are two distinct vertices. By the definition of $G$, the edge $xy$ is absent from $G$.
The number of induced copies of $F$ in $G$ that contain the non‑edge $xy$ is
$\frac{(m-1)_{r-1}k^{2r-\ell}\binom{k}{2}^{\ell-r-1}}{(\ell-r-1)!(2r-\ell)!}=\frac{(m-1)_{r-1}\bigl(k^{\ell-2}+o(k^{\ell-2})\bigr)}{2^{\ell-r-1}(\ell-r-1)!(2r-\ell)!}.$
If $\ell=2r-1$, then $F$ has no part of size larger than $1$ apart from the two vertices $x,y$; consequently, there is no induced copy of $F$ in $G\oplus xy$ that contains the edge $xy$. Otherwise, the number of such copies is
$\frac{(m-1)_{r-2}k^{2r-\ell-2}\binom{k}{2}^{\ell-r}}{(\ell-r)!(2r-\ell-2)!}=\frac{(m-1)_{r-2}\bigl(k^{\ell-2}+o(k^{\ell-2})\bigr)}{2^{\ell-r}(\ell-r)!(2r-\ell-2)!}.$
Therefore, regardless of the relation between $\ell$ and $r$, we have
$I(F,G)-I(F,G\oplus xy)$ $=\frac{(m-1)_{r-2}\bigl(2(\ell-r)(m-r+1)-(2r-\ell)(2r-\ell-1)\bigr)}{2^{\ell-r}(\ell-r)!(2r-\ell)!}\bigl(k^{\ell-2}+o(k^{\ell-2})\bigr)$
$=\frac{(m-1)_{r-2}\bigl(2(\ell-r)m+\ell-\ell^{2}+2r\ell-2r^{2}\bigr)}{2^{\ell-r}m^{\ell-2}(\ell-r)!(2r-\ell)!}\bigl(n^{\ell-2}+o(n^{\ell-2})\bigr).$
By Lemma 4.2 (i), $m>\frac{\ell(r-1)}{2(\ell-r)}$. Substituting this lower bound gives
$2(\ell-r)m+\ell-\ell^{2}+2r\ell-2r^{2}>\ell(r-1)+\ell-\ell^{2}+2r\ell-2r^{2}=(\ell-r)(2r-\ell).$
Since $\ell-r$ and $2r$-$\ell$ are positive integers, their product is at least 11. Hence for all sufficiently large $n$,
$I(F,G)-I(F,G\oplus xy)>\varepsilon n^{\ell-2}.$
This completes the proof of Claim 4.5. ∎
Combining Claim 4.4 and Claim 4.5, (S1) holds. Now we turn to verify (S2).
Claim 4.6.
For every set $A\subseteq[m]$ with $|A|\neq m-1$,
$I(F,G_{A^{*}})-I(F,G_{A})\geq\varepsilon n^{\ell-1},$
where $A^{*}$ is any subset of $[m]$ with $|A^{*}|=m-1$.
Proof.
For any subset $B\subseteq[m]$ with $|B|=q$, we proceed to compute $I(F,G_{B})-I(F,G)$. Note that the induced copies (say $T$) of $F$ in $G_{B}$ containing the new vertex $v_{B}$ fall into two categories:
Case 1: $v_{B}$ is adjacent to every other vertex of $T$. If q<r1q<r-1, the number of the induced copies of these $T$ is 0; otherwise this number is
$\frac{(q)_{r-1}\cdot k^{2r-\ell-1}\cdot\binom{k}{2}^{\ell-r}}{(\ell-r)!(2r-\ell-1)!}=\frac{(q)_{r-1}\bigl(k^{\ell-1}+o(k^{\ell-1})\bigr)}{2^{\ell-r}(\ell-r)!(2r-\ell-1)!}.$
Case 2: $v_{B}$ is not adjacent to every other vertex of $T$. If q<r1q<r-1, the number of the induced copies of these $T$ is 0; otherwise this number is
$\frac{(q)_{r-1}\cdot(m-q)\cdot k^{2r-\ell+1}\cdot\binom{k}{2}^{\ell-r-1}}{(\ell-r-1)!(2r-\ell)!}=\frac{(m-q)(q)_{r-1}\bigl(k^{\ell-1}+o(k^{\ell-1})\bigr)}{2^{\ell-r-1}(\ell-r-1)!(2r-\ell)!}.$
Adding these two cases together, regardless of the value of $q$, we obtain
$I(F,G_{B})-I(F,G)=\frac{(q)_{r-1}\cdot\bigl(2r-\ell+2(m-q)(\ell-r)\bigr)\cdot\bigl(k^{\ell-1}+o(k^{\ell-1})\bigr)}{2^{\ell-r}(\ell-r)!(2r-\ell)!}.$ (26)
Define for integers $q\in[r-1,m]$,
$h(q)\coloneqq(q)_{r-1}\cdot\bigl(2r-\ell+2(m-q)(\ell-r)\bigr).$ (27)
For convenience, we extend $h$ by setting $h(q)=0$ for all integers q<r1q<r-1. We will show that for all integers $q\in[m]\setminus\{m-1\}$,
$h(q)<h(m-1).$ (28)
Since $h(m-1)=\ell(m-1)_{r-1}$, it suffices to prove that for every integer $q\in[r-1,m]\setminus\{m-1\}$,
$(q)_{r-1}\cdot\bigl(2r-\ell+2(m-q)(\ell-r)\bigr)<\ell\cdot(m-1)_{r-1}.$ (29)
For $q=m$, inequality (29) becomes $(m)_{r-1}(2r-\ell)<\ell(m-1)_{r-1}$, which simplifies to $2(\ell-r)m>\ell(r-1)$. This is precisely the lower bound on $m$ given by Lemma 4.2 (i), and hence this case holds.
We now prove (29) by reverse induction on $q$, starting from q=m2q=m-2 down to q=r1q=r-1.
For the base case q=m2q=m-2, we need to show $(m-2)_{r-1}(3\ell-2r)<\ell(m-1)_{r-1}$, which is equivalent to $2(\ell-r)m<3r\ell-2r^{2}$-$\ell$. By the upper bound in Lemma 4.2 (ii), we have $m<\frac{\ell(\ell-1)}{2(\ell-r)}$; consequently,
$2(\ell-r)m-(3r\ell-2r^{2}-\ell)<\ell(\ell-1)-3r\ell+2r^{2}+\ell=(\ell-r)(\ell-2r)<0,$
where the last inequality uses $r<\ell<2r$. Thus the base case holds.
Now we turn to inductive step. Assume that (29) holds for some integer $q$ with $r\leq q\leq m-2$. We now prove it for q1q-1. Using the inductive hypothesis, we have
$\ \ell(m-1)_{r-1}-(q-1)_{r-1}\bigl(2r-\ell+2(m-q+1)(\ell-r)\bigr)$
$>$ $\ (q)_{r-1}\bigl(2r-\ell+2(m-q)(\ell-r)\bigr)-(q-1)_{r-1}\bigl(2r-\ell+2(m-q+1)(\ell-r)\bigr)$
$=$ $\ (q-1)_{r-2}\Bigl(2(\ell-r)\bigl((r-1)(m+1)-rq\bigr)+(r-1)(2r-\ell)\Bigr)$
$\geq$ $\ (q-1)_{r-2}\Bigl(2(\ell-r)(3r-m-1)+(r-1)(2r-\ell)\Bigr),$
where the last inequality follows from the assumption $q\leq m-2$. Indeed,
$(r-1)(m+1)-rq\geq(r-1)(m+1)-r(m-2)=3r-m-1.$
Using the upper bound $m<\frac{\ell(\ell-1)}{2(\ell-r)}$ from Lemma 4.2 (ii), we further have
$2(\ell-r)(3r-m-1)+(r-1)(2r-\ell)$ $>2(\ell-r)\left(3r-1-\frac{\ell(\ell-1)}{2(\ell-r)}\right)+(r-1)(2r-\ell)$
$=2(\ell-r)(3r-1)-\ell(\ell-1)+(r-1)(2r-\ell)$
$=5r\ell-\ell^{2}-4r^{2}=(\ell-r)(4r-\ell)>0.$
This proves that $\ell(m-1)_{r-1}-(q-1)_{r-1}\bigl(2r-\ell+2(m-q+1)(\ell-r)\bigr)>0,$ which is precisely (29) for q1q-1. Thus (28) holds.
Combining (26), (27), and (28), we obtain that for all sufficiently large $n$ and for any $A,A^{*}\subseteq[m]$ with $|A^{*}|=m-1$ and $|A|=p\neq m-1$,
$I(F,G_{A^{\ast}})-I(F,G_{A})$ $=\bigl(I(F,G_{A^{\ast}})-I(F,G)\bigr)-\bigl(I(F,G_{A})-I(F,G)\bigr)$
$=\bigl(h(m-1)-h(p)\bigr)\cdot\frac{k^{\ell-1}+o(k^{\ell-1})}{2^{\ell-r}(\ell-r)!(2r-\ell)!}\geq\frac{k^{\ell-1}+o(k^{\ell-1})}{2^{\ell-r}(\ell-r)!(2r-\ell)!}$
$=\frac{n^{\ell-1}+o(n^{\ell-1})}{2^{\ell-r}m^{\ell-1}(\ell-r)!(2r-\ell)!}>\varepsilon n^{\ell-1}.$
Hence, Claim 4.6 holds. ∎
By Theorem 4.1, all almost balanced complete $r$-partite graphs on $\ell\in[r+1,~2r-1]$ vertices is perfectly stable. This completes the proof of Theorem 1.5 for $\ell\leq 2r-1$. ∎
4.3 Proof of Theorem 1.5 for $\ell\geq 2r$
In this subsection, we prove Theorem 1.5 for the case $\ell\geq 2r$. The proof parallels that of the previous section, with several steps simplified by the fact that $\ell\geq 2r$ implies $a_{r}\geq 2$.
Proof of Theorem 1.5.
Let $F=K_{a_{1},\ldots,a_{r}}$ be an almost balanced complete $r$-partite graph on $\ell\geq 2r$ vertices, that is, $\ell=a_{1}+\cdots+a_{r}$. Let $m=m_{r,\ell}$ be the integer given by Lemma 3.2. It follows from Theorem 3.1 that $\mathrm{OPT}(F)=\{\bm{m}\}$ and
$i(F)=\kappa_{F}\cdot\frac{(m-1)_{r-1}}{m^{\ell-1}}.$
Assume that $a_{1}\geq\cdots\geq a_{r}$. If $a_{r}=1$, then the almost-balanced condition of $F$ forces $a_{1}\leq 2$, which implies $\ell\leq 2r-1$. This contradicts our assumption $\ell\geq 2r$. Therefore, we must have $a_{r}\geq 2$. Let $b_{1}>\cdots>b_{t}$ denote the distinct elements of the multiple set $\{\hskip-2.5pt\{\hskip 1.00006pta_{1},\ldots,a_{r}\hskip 1.00006pt\}\hskip-2.5pt\}$, occurring with multiplicities $r_{1},\ldots,r_{t}$, respectively. It follows from Lemma 4.2 (ii) that
$m<\frac{r\ell^{2}}{\ell^{2}-r^{2}}\leq\frac{r(2r)^{2}}{(2r)^{2}-r^{2}}=\frac{4r}{3}.$ (30)
Thus, by Theorem 4.1, it suffices to verify (S1) and (S2) with some constant $\varepsilon>0$ and all sufficiently large $n$.
Fix a sufficiently large $n$. For simplicity, we will use $o(\cdot)$ to denote lower-order terms and assume that $n=km$ for some integer $k$ (i.e., $n$ is divisible by $m$). It will be clear from the proof that this divisibility assumption has no substantive effect; it is included only to allow us to omit floor and ceiling signs.
Let $G\coloneqq G_{n,\bm{m}}$ be the $n$-vertex realization with parts $V_{1},\ldots,V_{m}$, where $|V_{i}|=k$ for all $i\in[m]$. Let $\varepsilon>0$ be the constant satisfying
$1/\varepsilon=m^{\ell+1}\cdot\mathrm{sym}(a_{1},\ldots,a_{r})\prod_{j=1}^{r}(a_{j})!.$
We first verify (S1) in the following two claims.
Claim 4.7.
Suppose that $\{x,y\}\subseteq V_{\alpha}$ for some $\alpha\in[m]$. Then
$I(F,G)-I(F,G\oplus xy)\geq\varepsilon n^{\ell-2}.$
Proof.
Every part of $F$ has size at least $2$, therefore an edge inside a part $V_{\alpha}$ cannot appear in any induced copy of $F$. Hence, for $x,y$ lying in the part $V_{\alpha}$ and for all sufficiently large $n$,
$I(F,G)-I(F,G\oplus xy)$ $=\sum_{i=1}^{t}\frac{(m-1)_{r-1}\binom{k-2}{b_{i}-2}\binom{k}{b_{i}}^{r_{i}-1}\prod_{j\in[t]\setminus\{i\}}\binom{k}{b_{j}}^{r_{j}}}{\mathrm{sym}(a_{1},\ldots,a_{r})/r_{i}}$
$=\frac{(m-1)_{r-1}\bigl(\sum_{i=1}^{r}a_{i}(a_{i}-1)\bigr)}{\mathrm{sym}(a_{1},\ldots,a_{r})\prod_{j=1}^{r}a_{j}!}\cdot\bigl(k^{\ell-2}+o(k^{\ell-2})\bigr)$
$=\frac{(m-1)_{r-1}\bigl(\sum_{i=1}^{r}a_{i}(a_{i}-1)\bigr)}{m^{\ell-2}\cdot\mathrm{sym}(a_{1},\ldots,a_{r})\prod_{j=1}^{r}a_{j}!}\cdot\bigl(n^{\ell-2}+o(n^{\ell-2})\bigr)>\varepsilon n^{\ell-2}.$
The first equality follows by classifying the cases according to the size of the part of $F$ that contains both $x$ and $y$. Here the sum runs over all distinct part sizes $b_{i}$ of $F$. For each fixed $i$, the factor $\binom{k-2}{b_{i}-2}$ chooses the remaining $b_{i}-2$ vertices from the part $V_{\alpha}$, the factor $\binom{k}{b_{i}}^{r_{i}-1}$ chooses vertices from the other parts of size $b_{i}$, and $\prod_{j\in[t]\setminus\{i\}}\binom{k}{b_{j}}^{r_{j}}$ handles the remaining parts. The denominator $\mathrm{sym}(a_{1},\dots,a_{r})/r_{i}$ accounts for the symmetries among parts of the same size. This proves Claim 4.7. ∎
Claim 4.8.
Suppose that $(x,y)\in V_{i}\times V_{j}$ for distinct $i,j\in[m]$. Then
$I(F,G)-I(F,G\oplus xy)\geq\varepsilon n^{\ell-2}.$
Proof.
Assume without loss of generality that $x\in V_{1}$ and $y\in V_{2}$. Since $x$ and $y$ lie in different parts, the edge $xy$ is present in $G$.
We first compute the number of induced copies of $F$ in $G$ that contain the edge $xy$, which equals
$\sum_{\{i,j\}\subseteq[t]}\frac{(m-2)_{r-2}\binom{k-1}{b_{i}-1}\binom{k-1}{b_{j}-1}\binom{k}{b_{i}}^{r_{i}-1}\binom{k}{b_{j}}^{r_{j}-1}\prod_{s\in[t]\setminus\{i,j\}}\binom{k}{b_{s}}^{r_{s}}}{\mathrm{sym}(a_{1},\ldots,a_{r})/(r_{i}r_{j})}$
$\qquad+\sum_{i\in[t],r_{i}>1}\frac{(m-2)_{r-2}\binom{k-1}{b_{i}-1}^{2}\binom{k}{b_{i}}^{r_{i}-2}\prod_{j\in[t]\setminus\{i\}}\binom{k}{b_{j}}^{r_{j}}}{\mathrm{sym}(a_{1},\ldots,a_{r})/(r_{i}(r_{i}-1))}$
$={}$ $\frac{(m-2)_{r-2}\left(\sum_{\{i,j\}\subseteq[r]}a_{i}a_{j}\right)}{\mathrm{sym}(a_{1},\ldots,a_{r})\prod_{j=1}^{r}(a_{j})!}\cdot\bigl(k^{\ell-2}+o(k^{\ell-2})\bigr)$
$={}$ $\frac{(m-2)_{r-2}\left(\sum_{\{i,j\}\subseteq[r]}a_{i}a_{j}\right)}{m^{\ell-2}\cdot\mathrm{sym}(a_{1},\ldots,a_{r})\prod_{j=1}^{r}(a_{j})!}\cdot(n^{\ell-2}+o(n^{\ell-2})).$ (31)
The two expressions in the leftmost side correspond to whether $x$ and $y$ lie in parts of $F$ of the same size or of different sizes. More precisely:
  • The first expression sums over distinct $i,j\in[t]$ to count copies of $F$ with $x$ in a part of size $b_{i}$ and $y$ in a part of size $b_{j}$. For a fixed pair $\{i,j\}$, the factor $\binom{k-1}{b_{i}-1}\binom{k-1}{b_{j}-1}$ chooses the remaining vertices from the parts containing $x$ and $y$, $\binom{k}{b_{i}}^{r_{i}-1}\binom{k}{b_{j}}^{r_{j}-1}$ selects vertices from the other parts of sizes $b_{i}$ and $b_{j}$, and the product $\prod_{s\notin\{i,j\}}\binom{k}{b_{s}}^{r_{s}}$ handle the remaining parts. The denominator $\mathrm{sym}(a_{1},\dots,a_{r})/(r_{i}r_{j})$ accounts for the symmetries among parts of the same size.
  • The second expression sums over $i\in[t]$ to count copies of $F$ with both $x,y$ in two distinct parts of size $b_{i}$ (clearly, $r_{i}\geq 2$). Here $\binom{k-1}{b_{i}-1}^{2}$ chooses the remaining vertices from the two particular parts of size $b_{i}$, $\binom{k}{b_{i}}^{r_{i}-2}$ selects vertices from the other $r_{i}-2$ parts of size $b_{i}$, and the remaining factors are analogous. The denominator $\mathrm{sym}(a_{1},\dots,a_{r})/(r_{i}(r_{i}-1))$ accounts for the symmetries among parts of the same size correspondingly.
Suppose that $a_{r}\geq 3$. Since every part of $F$ has size at least $3$, the non‑edge $xy$ in $G\oplus xy$ cannot belong to any induced copy of $F$. Hence for sufficiently large $n$,
$I(F,G)-I(F,G\oplus xy)=\frac{(m-2)_{r-2}\left(\sum_{\{i,j\}\subseteq[r]}a_{i}a_{j}\right)}{m^{\ell-2}\cdot\mathrm{sym}(a_{1},\ldots,a_{r})\prod_{j=1}^{r}(a_{j})!}\cdot\bigl(n^{\ell-2}+o(n^{\ell-2})\bigr)>\varepsilon n^{\ell-2}.$
It remains to consider the case $a_{r}=2$. Note that in this case, $b_{t}=2$. So $r_{t}$ is the multiplicity of the part size $2$. Then the induced copies of $F$ in $G\oplus xy$ that contain the non-edge $xy$ correspond bijectively to the induced copies of $K_{a_{1},\ldots,a_{r-1}}$ in $G-V_{1}-V_{2}$ (i.e., the graph obtained from $G$ by deleting all vertices of $V_{1}$ and $V_{2}$). Thus the number of these copies is
$\frac{(m-2)_{r-1}}{\mathrm{sym}(a_{1},\ldots,a_{r-1})}\prod_{i=1}^{r-1}\binom{k}{a_{i}}=\frac{2(m-2)_{r-1}\cdot r_{t}}{\mathrm{sym}(a_{1},\ldots,a_{r})\prod_{j=1}^{r}(a_{j})!}\bigl(k^{\ell-2}+o(k^{\ell-2})\bigr).$ (32)
Combining (4.3)and (32), we obtain
$I(F,G)-I(F,G\oplus xy)=\frac{(m-2)_{r-2}\left(\big(\sum_{\{i,j\}\subseteq[r]}a_{i}a_{j}\big)-2(m-r)r_{t}\right)}{\mathrm{sym}(a_{1},\ldots,a_{r})\prod_{j=1}^{r}(a_{j})!}\cdot\bigl(k^{\ell-2}+o(k^{\ell-2})\bigr).$
Note that $a_{i}\geq 2$ for all $i$, so $\sum_{\{i,j\}\subseteq[r]}a_{i}a_{j}\geq\sum_{\{i,j\}\subseteq[r]}2\cdot 2=4\binom{r}{2}=4r(r-1)$. Using the bound $m<\frac{4r}{3}$ from (30), we obtain
$\sum_{\{i,j\}\subseteq[r]}a_{i}a_{j}-2r_{t}(m-r)$ $\geq 4r(r-1)-2r(m-r)=2r\bigl(3r-2-m\bigr)$
$>2r\Bigl(3r-2-\frac{4r}{3}\Bigr)=2r\cdot\frac{5r-6}{3}>0\qquad(\text{since }r\geq 2).$
Therefore, $I(F,G)-I(F,G\oplus xy)>\varepsilon n^{\ell-2}$ for all sufficiently large $n$. This proves Claim 4.8. ∎
Combining Claim 4.7 and Claim 4.8, we derive that (S1) holds. Next we verify (S2).
Claim 4.9.
For every set $A\subseteq[m]$ with $|A|\neq m-1$,
$I(F,G_{A^{*}})-I(F,G_{A})\geq\varepsilon n^{\ell-1},$
where $A^{*}$ is any subset of $[m]$ with $|A^{*}|=m-1$.
Proof.
For any subset $B\subseteq[m]$ with $|B|=q$, we proceed to compute $I(F,G_{B})-I(F,G)$.
If $q=m$ or q<r1q<r-1, no induced copy of $F$ can contain $v_{B}$ (in the former case $v_{B}$ is adjacent to all parts, contradicting that $a_{r}\geq 2$; in the latter case $v_{B}$ is adjacent to fewer than r1r-1 parts, while every vertex of $F$ is adjacent to vertices from r1r-1 other parts). Otherwise, the number of such copies is given by
$\sum_{i=1}^{t}\frac{(m-q)(q)_{r-1}\binom{k}{b_{i}-1}\binom{k}{b_{i}}^{r_{i}-1}\prod_{j\in[t]\setminus\{i\}}\binom{k}{b_{j}}^{r_{j}}}{\mathrm{sym}(a_{1},\ldots,a_{r})/r_{i}}=\frac{\ell(m-q)(q)_{r-1}}{\mathrm{sym}(a_{1},\ldots,a_{r})\prod_{j=1}^{r}a_{j}!}\bigl(k^{\ell-1}+o(k^{\ell-1})\bigr),$
where the summation enumerates all placements of $v_{B}$ according to the size $b_{i}$ of the part of $F$ in which it lies, and we use the identity $\sum_{i}r_{i}b_{i}=\ell$. Hence we obtain
$I(F,G_{B})-I(F,G)=\frac{\ell(m-q)(q)_{r-1}}{\mathrm{sym}(a_{1},\ldots,a_{r})\prod_{j=1}^{r}a_{j}!}\bigl(k^{\ell-1}+o(k^{\ell-1})\bigr).$ (33)
Define for integers $q\in[r-1,m-1]$
$H(q)\coloneqq(m-q)(q)_{r-1}.$ (34)
For convenience, we set $H(q)=0$ for all integers $q\leq r-2$. We will show that for every integer q<m1q<m-1,
$H(q)<H(m-1).$ (35)
Since $H(m-1)=(m-1)_{r-1}$, it suffices to prove that for every $q\in[r-1,m-2]$,
$(m-1)_{r-1}-(m-q)(q)_{r-1}>0.$ (36)
We prove (36) by reverse induction on $q$. For the base case q=m2q=m-2,
$(m-1)_{r-1}-2(m-2)_{r-1}=\frac{2r-1-m}{m-r}\,(m-2)_{r-1}.$
From (30) we have $m<4r/3$ and thus $2r-1-m>2r-1-4r/3=(2r-3)/3\geq 1/3$ (since $r\geq 2$). Hence the expression is positive. Now we turn to inductive step. Assume (36) holds for some $q$ with $r\leq q\leq m-2$. We now prove it for q1q-1, Using the inductive hypothesis, we have
$(m-1)_{r-1}-(m-q+1)(q-1)_{r-1}>(m-q)(q)_{r-1}-(m-q+1)(q-1)_{r-1}$
$={}$ $(q-1)_{r-1}\,\frac{(r-1)(m+1)-rq}{\,q-r+1\,}\geq(q-1)_{r-1}\,\frac{3r-m-1}{\,q-r+1\,}$,
where the last inequality follows from the assumption $q\leq m-2$. Using again $m<4r/3$ gives 3rm1>3r4r/31=(5r3)/3>03r-m-1>3r-4r/3-1=(5r-3)/3>0 for $r\geq 2$. Therefore
$(m-1)_{r-1}-(m-q+1)(q-1)_{r-1}>0,$
which is precisely (36) for q1q-1. By induction, we obtain (35).
Combining (33), (34) and (35), we obtain for all sufficiently large $n$ and any $A,A^{*}\subseteq[m]$ with $|A^{*}|=m-1$, $|A|=p\neq m-1$,
$I(F,G_{A^{*}})-I(F,G_{A})$ $=\bigl(I(F,G_{A^{*}})-I(F,G)\bigr)-\bigl(I(F,G_{A})-I(F,G)\bigr)$
$=\bigl(H(m-1)-H(p)\bigr)\cdot\frac{k^{\ell-1}+o(k^{\ell-1})}{\mathrm{sym}(a_{1},\ldots,a_{r})\prod_{j=1}^{r}a_{j}!}\geq\frac{k^{\ell-1}+o(k^{\ell-1})}{\mathrm{sym}(a_{1},\ldots,a_{r})\prod_{j=1}^{r}a_{j}!}$
$=\frac{n^{\ell-1}+o(n^{\ell-1})}{m^{\ell-1}\cdot\mathrm{sym}(a_{1},\ldots,a_{r})\prod_{j=1}^{r}a_{j}!}>\varepsilon n^{\ell-1}.$
Hence Claim 4.9 holds. ∎
Therefore, by Theorem 4.1, the inducibility for $F$ is perfectly stable, completing Theorem 1.5. ∎
As a corollary of Theorem 1.5, for any almost balanced complete $r$-partite graph $F$ and sufficiently large $n$, all extremal graphs for $I(F,n)$ are complete multipartite. In Section 5, we strengthen this by showing that the extremal graph for $I(F,n)$ is unique, namely, the Turán graph $T_{m}(n)$ for $m=m_{r,\ell}$.
5 Exact results
In this section, we show that extremal graphs for $I(F,n)$ are Turán graphs whenever $F$ is an almost balanced complete multipartite graph, thus proving Theorem 1.1. First, we prove a technical lemma.
Lemma 5.1.
Let $s,t\geq 2$ be integers. Suppose $x,y,N$ are real numbers with $y>x+1$ satisfying |xN|=o(N)|x-N|=o(N) and |yN|=o(N)|y-N|=o(N) as $N\to\infty$. Then
$\tbinom{x+1}{s}+\tbinom{y-1}{s}-\tbinom{x}{s}-\tbinom{y}{s}=-\tfrac{1}{(s-2)!}(y-x-1+o(1))N^{s-2},$
and
$\tbinom{x+1}{t}\tbinom{y-1}{s}+\tbinom{x+1}{s}\tbinom{y-1}{t}-\tbinom{x}{t}\tbinom{y}{s}-\tbinom{x}{s}\tbinom{y}{t}=\tfrac{t+s-(t-s)^{2}}{s!t!}\big(y-x-1+o(1)\big)N^{t+s-2}.$
Proof.
By the Mean Value Theorem, there exists $z_{0}$ between $x$ and y1y-1 (hence $|z_{0}-N|=o(N)$) such that
$\tbinom{x+1}{s}+\tbinom{y-1}{s}-\tbinom{x}{s}-\tbinom{y}{s}$ $=\tbinom{x}{s-1}-\tbinom{y-1}{s-1}=(x-y+1)\bigl(\tbinom{z}{s-1}\bigr)^{\prime}\big|_{z=z_{0}}$
$=\tfrac{1}{(s-2)!}(x-y+1)\bigl(z_{0}^{s-2}+o(z_{0}^{s-2})\bigr)=-\tfrac{1}{(s-2)!}\bigl(y-x-1+o(1)\bigr)N^{s-2}.$
For the second formula, let
$g(x,y)\coloneqq(x+1)(y-s)_{t-s+1}+(x+1)(y-s)(x-s+1)_{t-s}-(x-s+1)y(y-s)_{t-s}-y(x-s+1)_{t-s+1}.$
A direct computation gives
$g(x,y)=s\bigl(y(y-s)_{t-s}-(x+1)(x+1-s)_{t-s}\bigr)-t(x+1)y\Bigl(\tfrac{(y-s)_{t-s}}{y}-\tfrac{(x+1-s)_{t-s}}{x+1}\Bigr).$
Applying the Mean Value Theorem to the two differences, we find $z_{1},z_{2}$ between $x$ and y1y-1 with $|z_{i}-N|=o(N)$ such that
$y(y-s)_{t-s}-(x+1)(x+1-s)_{t-s}$ $=(y-x-1)\bigl(z(z-s)_{t-s}\bigr)^{\prime}\big|_{z=z_{1}}$
$=\bigl(y-x-1+o(1)\bigr)(t-s+1)N^{t-s},$
$\tfrac{(y-s)_{t-s}}{y}-\tfrac{(x+1-s)_{t-s}}{x+1}$ $=(y-x-1)\Bigl(\tfrac{(z-s)_{t-s}}{z}\Bigr)^{\prime}\big|_{z=z_{2}}$
$=\bigl(y-x-1+o(1)\bigr)(t-s-1)N^{t-s-2}.$
Since $x,y=N+o(N)$, we have $t(x+1)y=tN^{2}+o(N^{2})$. Substituting these asymptotics into $g(x,y)$ yields
$g(x,y)=\bigl(y-x-1+o(1)\bigr)\bigl(t+s-(t-s)^{2}\bigr)N^{t-s}.$
Finally, using $(x)_{s-1}=N^{s-1}+o(N^{s-1})$ and $(y-1)_{s-1}=N^{s-1}+o(N^{s-1})$,
$\tbinom{x+1}{t}\tbinom{y-1}{s}+\tbinom{x+1}{s}\tbinom{y-1}{t}-\tbinom{x}{t}\tbinom{y}{s}-\tbinom{x}{s}\tbinom{y}{t}$
$\quad=\frac{(x)_{s-1}(y-1)_{s-1}}{t!\,s!}\,g(x,y)=\frac{t+s-(t-s)^{2}}{s!\,t!}\,\bigl(y-x-1+o(1)\bigr)N^{t+s-2}.$
This completes the proof. ∎
We are now ready to prove Theorem 1.1.
Proof of Theorem 1.1.
Let $F$ be an almost balanced complete $r$-partite graph on $\ell$ vertices with part sizes $a_{1}\geq\cdots\geq a_{r}$. Let $n$ be a sufficiently large integer and set $m\coloneqq m_{r,\ell}$. By Theorem 3.1 and Theorem 1.5, every $n$-vertex graph $G$ attaining $I(F,n)$ is a complete $m$-partite graph whose parts have sizes $(1+o(1))N$, where $N\coloneqq n/m$.
We now determine the exact part sizes. Write the part sizes of $G$ as $n_{1}\geq n_{2}\geq\dots\geq n_{m}$, with $\sum_{i=1}^{m}n_{i}=n$ and $|n_{i}-N|=o(n)$ for all $i$. Let $y\coloneqq n_{1}$ and $x\coloneqq n_{m}$; by definition $y\geq x$.
Claim 5.2.
$\binom{\ell}{2}>m\sum_{k=1}^{r}\binom{a_{k}}{2}$.
Proof.
Since $F$ is almost balanced, we have $(a_{i}-a_{j})^{2}\leq a_{i}+a_{j}-2$ for all $1\leq i<j\leq r$. Therefore,
$r\sum_{k=1}^{r}{a_{k}}^{2}=\Bigl(\sum_{k=1}^{r}a_{k}\Bigr)^{2}+\sum_{1\leq i<j\leq r}(a_{i}-a_{j})^{2}\leq\ell^{2}+\sum_{1\leq i<j\leq r}(a_{i}+a_{j}-2)=\ell^{2}+(r-1)\ell-r(r-1).$
Consequently,
$r\sum_{k=1}^{r}\binom{a_{k}}{2}\leq\binom{\ell}{2}-\binom{r}{2}$. (37)
By Lemma 4.2 (ii), we have
$m<\frac{r\ell^{2}}{\ell^{2}-r^{2}}=\frac{r\ell^{2}}{(\ell-r)(\ell+r)}<\frac{r\ell(\ell-1)}{(\ell-r-1)(\ell+r)}=\frac{r\binom{\ell}{2}}{\binom{\ell}{2}-\binom{r}{2}}$. (38)
Combining (37) and (38) yields $\binom{\ell}{2}>m\sum_{k=1}^{r}\binom{a_{k}}{2}$. ∎
For $\{i,j\}\subseteq[r]$, define
$A_{i}$ $\coloneqq\sum_{(s_{k})_{k\in[r]\setminus\{i\}}}\prod_{k\in[r]\setminus\{i\}}\tbinom{n_{s_{k}}}{a_{k}}=\left(\frac{(m-2)_{r-1}\cdot a_{i}!}{\prod_{k\in[r]}a_{k}!}+o(1)\right)N^{\ell-a_{i}},$
$B_{i,j}$ $\coloneqq\sum_{(s_{k})_{k\in[r]\setminus\{i,j\}}}\prod_{k\in[r]\setminus\{i,j\}}\tbinom{n_{s_{k}}}{a_{k}}=\left(\frac{(m-2)_{r-2}\cdot a_{i}!a_{j}!}{\prod_{k\in[r]}a_{k}!}+o(1)\right)N^{\ell-a_{i}-a_{j}},$
where the tuples $(s_{k})$ run over $\big([2,m-1]\big)_{r-1}$ for $A_{i}$ and run over $\big([2,m-1]\big)_{r-2}$ for $B_{i,j}$ (the indices $1$ and $m$ correspond to the parts of sizes $y$ and $x$, which will be treated separately).
Set $\hat{I}(F,H)\coloneqq\mathrm{sym}(a_{1},\dots,a_{r})\cdot I(F,H)$ for any graph $H$. Let $G_{*}\cong K_{n_{1}-1,n_{2},\dots,n_{m}+1}$. By (4),
$\hat{I}(F,G_{*})-\hat{I}(F,G)$ $=\sum_{i\in[r]}\left(\tbinom{x+1}{a_{i}}+\tbinom{y-1}{a_{i}}-\tbinom{x}{a_{i}}-\tbinom{y}{a_{i}}\right)A_{i}$
$\quad+\sum_{\{i,j\}\subseteq[r]}\left(\tbinom{x+1}{a_{i}}\tbinom{y-1}{a_{j}}+\tbinom{x+1}{a_{j}}\tbinom{y-1}{a_{i}}-\tbinom{x}{a_{i}}\tbinom{y}{a_{j}}-\tbinom{x}{a_{j}}\tbinom{y}{a_{i}}\right)B_{i,j}.$ (39)
If $a_{i}=1$, the coefficient of $A_{i}$ vanishes. Otherwise $a_{i}\geq 2$, Lemma 5.1 gives
$\left(\tbinom{x+1}{a_{i}}+\tbinom{y-1}{a_{i}}-\tbinom{x}{a_{i}}-\tbinom{y}{a_{i}}\right)A_{i}$
$={}$ $-\frac{1}{(a_{i}-2)!}\bigl(y-x-1+o(1)\bigr)N^{a_{i}-2}\cdot\left(\frac{(m-2)_{r-1}\cdot a_{i}!}{\prod_{k\in[r]}a_{k}!}+o(1)\right)N^{\ell-a_{i}}$
$={}$ $-\frac{\bigl(y-x-1+o(1)\bigr)(m-2)_{r-1}}{\prod_{k\in[r]}a_{k}!}\cdot a_{i}(a_{i}-1)\,N^{\ell-2}.$ (40)
(Note that when $a_{1}=1$, (5) holds trivially because both sides are zero.) Similarly, Lemma 5.1 yields
$\left(\tbinom{x+1}{a_{i}}\tbinom{y-1}{a_{j}}+\tbinom{x+1}{a_{j}}\tbinom{y-1}{a_{i}}-\tbinom{x}{a_{i}}\tbinom{y}{a_{j}}-\tbinom{x}{a_{j}}\tbinom{y}{a_{i}}\right)B_{i,j}$
$={}$ $\frac{a_{i}+a_{j}-(a_{i}-a_{j})^{2}}{a_{i}!a_{j}!}\bigl(y-x-1+o(1)\bigr)N^{a_{i}+a_{j}-2}\cdot\left(\frac{(m-2)_{r-2}\cdot a_{i}!a_{j}!}{\prod_{k\in[r]}a_{k}!}+o(1)\right)N^{\ell-a_{i}-a_{j}}$
$={}$ $\frac{\bigl(y-x-1+o(1)\bigr)(m-2)_{r-2}}{\prod_{k\in[r]}a_{k}!}\cdot\bigl(a_{i}+a_{j}-(a_{i}-a_{j})^{2}\bigr)\,N^{\ell-2}.$ (41)
Substituting (5) and (5) into (5) gives
$\hat{I}(F,G_{*})-\hat{I}(F,G)={}$ $\frac{\bigl(y-x-1+o(1)\bigr)(m-2)_{r-2}}{\prod_{k\in[r]}a_{k}!}N^{\ell-2}$
$\times\Biggl(-(m-r)\sum_{i\in[r]}a_{i}(a_{i}-1)+\sum_{\{i,j\}\subseteq[r]}\bigl(a_{i}+a_{j}-(a_{i}-a_{j})^{2}\bigr)\Biggr).$
Straightforward calculations show that
$-(m-r)\sum_{i\in[r]}a_{i}(a_{i}-1)+\sum_{\{i,j\}\subseteq[r]}\big(a_{i}+a_{j}-(a_{i}-a_{j})^{2}\big)=2\Bigl(\tbinom{\ell}{2}-m\sum_{i\in[r]}\tbinom{a_{i}}{2}\Bigr)$.
Therefore,
$\hat{I}(F,G_{*})-\hat{I}(F,G)=\frac{2\bigl(y-x-1+o(1)\bigr)(m-2)_{r-2}}{\prod_{k\in[r]}a_{k}!}\cdot N^{\ell-2}\Bigl(\tbinom{\ell}{2}-m\sum_{i\in[r]}\tbinom{a_{i}}{2}\Bigr)$. (42)
By Claim 5.2, the factor $\binom{\ell}{2}-m\sum_{i}\binom{a_{i}}{2}$ is positive. If $y>x+1$, then for sufficiently large $n$ the term yx1+o(1)y-x-1+o(1) is also positive, implying $\hat{I}(F,G_{*})>\hat{I}(F,G)$. This shows that $I(F,G_{*})>I(F,G)$ as well, contradicting the maximality of $G$. Hence we must have $y\leq x+1$, which forces all part sizes of $G$ to differ by at most $1$, i.e., $G$ is a Turán graph $T_{m}(n)$. This completes the proof of Theorem 1.1. ∎
For the balanced case $K_{r}(t)=T_{r}(rt)$, the requirement that $n$ be sufficiently large can be dropped, though the extremal graph may no longer be unique.
Theorem 5.3.
Let $r,t\geq 2$ be integers. For every integer $n$, there exists an integer $k$ such that the Turán graph $T_{k}(n)$ is an extremal graph for $I(K_{r}(t),n)$.
Proof.
The case $n<rt$ is trivial. For $n\geq rt$, Lemma 2.2 guarantees that some complete multipartite graph $G$ is extremal for $I(K_{r}(t),n)$. Let $G$ have $k$ parts with sizes $a_{1}\geq\dots\geq a_{k}$. Since $G$ must accommodate an induced copy of $K_{r}(t)$, we necessarily have $k\geq r$ and $a_{k}\geq t$.
Set $F\coloneqq K_{r}(t)$. Define $G^{\prime}$ as the complete $k$-partite graph with part sizes $a_{1}-1,a_{2},\dots,a_{k-1},a_{k}+1$, and $G^{\prime\prime}$ as the complete (k1)(k-1)-partite graph with part sizes $a_{1}+a_{k},a_{2},\dots,a_{k-1}$.
Claim 5.4.
If $a_{1}-a_{k}>1$, then $I(F,G)<\max\{I(F,G^{\prime}),I(F,G^{\prime\prime})\}$.
Proof.
Let $x\coloneqq a_{k}$ and $y\coloneqq a_{1}$. Then $y-1>x\geq t$. Suppose for a contradiction that both $I(F,G)\geq I(F,G^{\prime})$ and $I(F,G)\geq I(F,G^{\prime\prime})$ hold. Define
$A\coloneqq r\sum_{(i_{1},\dots,i_{r-1})\in([k-1]\setminus\{1\})_{r-1}}\tbinom{a_{i_{1}}}{t}\cdots\tbinom{a_{i_{r-1}}}{t},\quad B\coloneqq r(r-1)\sum_{(i_{1},\dots,i_{r-2})\in([k-1]\setminus\{1\})_{r-2}}\tbinom{a_{i_{1}}}{t}\cdots\tbinom{a_{i_{r-2}}}{t}$.
Since $a_{i}\geq t$ for all $i$ and $k\geq r\geq 2$, the sum defining $B$ contains at least one term (as $k-1\geq r-2$), and each term is positive because $\binom{a_{i_{j}}}{t}\geq 1$; therefore $B>0$. By (4), we obtain the two inequalities
$0$ $\leq r!\bigl(I(F,G)-I(F,G^{\prime})\bigr)=\alpha_{1}A-\beta_{1}B,$ (43)
$0$ $\leq r!\bigl(I(F,G)-I(F,G^{\prime\prime})\bigr)=\beta_{2}B-\alpha_{2}A,$ (44)
where
$\alpha_{1}=\tbinom{x}{t}+\tbinom{y}{t}-\tbinom{x+1}{t}-\tbinom{y-1}{t},~\alpha_{2}=\tbinom{x+y}{t}-\tbinom{x}{t}-\tbinom{y}{t},~\beta_{1}=\tbinom{x+1}{t}\tbinom{y-1}{t}-\tbinom{x}{t}\tbinom{y}{t},~\beta_{2}=\tbinom{x}{t}\tbinom{y}{t}$.
Observe that since $y>x\geq t\geq 2$, we have
$\alpha_{1}=\tbinom{y-1}{t-1}-\tbinom{x}{t-1}>0,\quad\alpha_{2}=\sum_{i=1}^{t-1}\tbinom{x}{i}\tbinom{y}{t-i}>0,\quad\beta_{1}=\tfrac{y-x-1}{t}\tbinom{x}{t-1}\tbinom{y-1}{t-1}>0,\quad\beta_{2}>0.$
Combining inequalities (43) and (44) and using $B>0$, we obtain the necessary condition
$\frac{\beta_{1}}{\alpha_{1}}\leq\frac{A}{B}\leq\frac{\beta_{2}}{\alpha_{2}},$
which, after rearrangement, is equivalent to
$0\leq\beta_{2}\alpha_{1}-\beta_{1}\alpha_{2}$ $=\tbinom{x}{t}\tbinom{y}{t}\Bigl(\tbinom{x}{t}+\tbinom{y}{t}-\tbinom{x+1}{t}-\tbinom{y-1}{t}\Bigr)-\Bigl(\tbinom{x+1}{t}\tbinom{y-1}{t}-\tbinom{x}{t}\tbinom{y}{t}\Bigr)\Bigl(\tbinom{x+y}{t}-\tbinom{x}{t}-\tbinom{y}{t}\Bigr)$
$=\tbinom{y}{t}\tbinom{y-1}{t}\Bigl(\tbinom{x+1}{t}-\tbinom{x}{t}\Bigr)-\tbinom{x+1}{t}\tbinom{x}{t}\Bigl(\tbinom{y}{t}-\tbinom{y-1}{t}\Bigr)-\tbinom{x+y}{t}\Bigl(\tbinom{x+1}{t}\tbinom{y-1}{t}-\tbinom{x}{t}\tbinom{y}{t}\Bigr)$
$=\tbinom{y}{t}\tbinom{y-1}{t}\tbinom{x}{t-1}-\tbinom{x+1}{t}\tbinom{x}{t}\tbinom{y-1}{t-1}-\tfrac{y-x-1}{t}\tbinom{x}{t-1}\tbinom{y-1}{t-1}\tbinom{x+y}{t}$
$=\tfrac{1}{t}\tbinom{x}{t-1}\tbinom{y-1}{t-1}\Bigl(y\tbinom{y-1}{t}-(x+1)\tbinom{x}{t}-(y-x-1)\tbinom{x+y}{t}\Bigr)$.
Consequently, we obtain
$y\tbinom{y-1}{t}-(x+1)\tbinom{x}{t}\geq(y-x-1)\tbinom{x+y}{t}$.
However, since $z\mapsto\binom{z}{t}$ is strictly convex on $t,\infty)$ for $t\geq 2$, Jensen’s inequality together with the convex combination $\frac{x+1}{y}\cdot x+\frac{y-x-1}{y}\cdot(x+y)=y-1$ gives
$\tfrac{x+1}{y}\tbinom{x}{t}+\tfrac{y-x-1}{y}\tbinom{x+y}{t}>\tbinom{y-1}{t}$,
which is equivalent to $y\tbinom{y-1}{t}-(x+1)\tbinom{x}{t}<(y-x-1)\tbinom{x+y}{t}$, contradicting the previous inequality. This completes the proof of Claim 5.4. ∎
By Claim 5.4, if $a_{1}-a_{k}>1$, then $G$ cannot be extremal. Therefore, in any extremal complete multipartite graph we must have $a_{1}-a_{k}\leq 1$; i.e., all part sizes differ by at most 11. Consequently, every extremal graph is isomorphic to a Turán graph $T_{k}(n)$. This completes the proof of Theorem 5.3. ∎
6 Concluding remarks
Let $F$ be a complete $r$-graph on $\ell\geq r+1$ vertices. Observe that if a maximizer $\bm{x}=(x_{1},x_{2},\ldots)\in\mathrm{OPT}(F)$ satisfies $x_{0}=0$ and $x_{i}=x_{j}$ for all ${i,j}\subseteq\mathrm{supp}(\bm{x})$, then necessarily $\bm{x}=\bm{m}_{r,\ell}$, where $m_{r,\ell}$ is the integer given by Lemma 3.2.
A natural next step beyond Theorem 1.1 is to understand to what extent the “almost balanced” hypothesis can be relaxed.
Problem 6.1.
For all $\ell>r\geq 3$, characterize the complete $r$-partite graphs $F$ on $\ell$ vertices for which $\mathrm{OPT}(F)=\{\bm{m}_{r,\ell}\}$.
Assume that $F=K_{a_{1},\ldots,a_{r}}$ is a complete $r$-partite graphs $F$ on $\ell$ vertices. Then a necessary condition for $F$ satisfying $\mathrm{OPT}(F)=\{\bm{m}_{r,\ell}\}$ is
$\binom{\ell}{2}\geq m_{r,\ell}\sum_{k=1}^{r}\binom{a_{k}}{2}$. (45)
Indeed, let $m=m_{r,\ell}$. If (45) fails, consider the perturbation (for a sufficiently small constant $\delta>0$)
$\bm{x}_{\delta}\coloneqq\big(\tfrac{1+\delta}{m},\underbrace{\tfrac{1}{m},\ldots,\tfrac{1}{m}}_{m-2},\tfrac{1-\delta}{m},0,\ldots\big).$
By (7), we have
$p_{F}(\bm{x}_{\delta})=\frac{\kappa_{F}(m-2)_{r-2}}{m^{\ell}}\bigl((m-r)(m-r-1)+(m-r)S(\delta)+T(\delta)\bigr),$
where
$S(\delta)\coloneqq\sum_{i\in[r]}\bigl((1+\delta)^{a_{i}}+(1-\delta)^{a_{i}}\bigr)\quad\text{and}\quad T(\delta)\coloneqq\sum_{\{i,j\}\subseteq[r]}(1+\delta)^{a_{i}}(1-\delta)^{a_{j}}.$
Writing $q(\delta)\coloneqq(m-r)S(\delta)+T(\delta)$, a direct calculation yields
$q^{\prime}(0)=0\quad\text{and}\quad q^{\prime\prime}(0)=2\Bigl(m\sum_{i=1}^{r}\tbinom{a_{i}}{2}-\tbinom{\ell}{2}\Bigr)>0.$
It follows that $\bm{m}$ cannot be a maximizer, as claimed.
It should be pointed out that (45) is not a sufficient condition for $\mathrm{OPT}(F)=\{\bm{m}_{r,\ell}\}$ to hold (even when replaced by a strict inequality). In forthcoming work [13], we will determine $i(F)$ and establish the perfect stability for all complete $3$-partite graphs $F$. In particular, our result will show that $F=K_{12,7,7}$ satisfies (45) with strict inequality. However, the unique maximizer in $\mathrm{OPT}(F)$ is $\left(\alpha,\frac{1-\alpha}{2},\frac{1-\alpha}{2},0,\ldots\right)$, where $\alpha\approx 0.396884$ is the largest real root of
$130x^{5}+25x^{4}-90x^{3}+80x^{2}-40x+7=0.$
On the other hand, the sufficient condition (1) is not necessary in general. For example, $F=K_{4,8,8}$ is not almost balanced (since $\binom{8-4}{2}=6>4$), yet our result will show that the unique maximizer for $F$ is $(1/3,1/3,1/3,0,\ldots)$.
Thus, the answer to Problem 6.1, in terms of the algebraic relations among $a_{1},\ldots,a_{r}$, lies strictly between the hypersurfaces defined by (1) and (45), and appears to be quite complicated.
Acknowledgments
X.L. was supported by the Excellent Young Talents Program (Overseas) of the National Natural Science Foundation of China. J.M. was supported by National Key Research and Development Program of China 2023YFA1010201 and National Natural Science Foundation of China grant 12125106. T.Z. was supported by Innovation Program for Quantum Science and Technology 2021ZD0302902. X.L. would like to thank Oleg Pikhurko for clarifying the necessity of considering the case $x_{0}>0$ in the proof of Theorem 3.1.